### v1

# SkyScenes: A Synthetic Dataset for Aerial Scene Understanding

Abstract
Real-world aerial scene understanding is limited by a lack of datasets that contain densely annotated images curated under a diverse set of conditions. Due to inherent challenges in obtaining such images in controlled real-world settings, we present SkyScenes, a synthetic dataset of densely annotated aerial images captured from Unmanned Aerial Vehicle (UAV) perspectives. We carefully curate SkyScenes images from Carla to comprehensively capture diversity across layout (urban and rural maps), weather conditions, times of day, pitch angles and altitudes with corresponding semantic, instance and depth annotations. Through our experiments using SkyScenes, we show that (1) Models trained on SkyScenes generalize well to different real-world scenarios, (2) augmenting training on real images with SkyScenes data can improve real-world performance, (3) controlled variations in SkyScenes can offer insights into how models respond to changes in viewpoint conditions, and (4) incorporating additional sensor modalities (depth) can improve aerial scene understanding.
| Dataset | Reproducibility | Diversity | Annotations | Image Capture | Scale | |||||||
| Metadata | Contr. Var. | Town | Daytime | Weather | Semantic | Instance | Depth | Altitude | Perspective | Resolution | ||
| Real | ||||||||||||
| 1 UAVid [19] | ✗ | ✗ | ✓ | ✗ | ✗ | ✓ | ✗ | ✗ | Med | Obl. | K | |
| 2 AeroScapes [22] | ✗ | ✗ | ✗ | ✗ | ✗ | ✓ | ✗ | ✗ | (Low, Med) | (Obl., Nad.) | K | |
| 3 ICG Drone [14] | ✗ | ✗ | ✗ | ✗ | ✗ | ✓ | ✗ | ✗ | Low | Nad. | K | |
| Synthetic | ||||||||||||
| 4 MidAir [9] | ✓ | partial | ✓ | ✓ | ✓ | ✓ | ✗ | ✓ | Low | (Obl., Nad.) | K | |
| 5 VALID [2] | ✗ | ✗ | ✓ | ✓ | ✓ | ✓ | ✓ | ✓ | (Low, Med, High) | Nad. | K | |
| 6 Espada [17] | ✗ | ✗ | ✓ | ✗ | ✗ | ✗ | ✗ | ✓ | (Med, High) | Nad. | K | |
| 7 TartanAir [35] | ✓ | * | ✓ | ✓ | ✓ | ✓ | ✗ | ✓ | Low | (Fwd., Obl.) | M | |
| 8 UrbanScene3D [16] | * | * | ✓ | * | * | ✗ | ✗ | ✗ | Med | Obl. | K | |
| 9 SynthAer [30] | ✓ | ✓ | ✗ | ✓ | ✗ | ✓ | ✗ | ✗ | (Low, Med) | Obl. | K | |
| 10 SynDrone [25] | ✓ | ✗ | ✓ | ✗ | ✗ | ✓ | ✗ | ✓ | (Low, Med, High) | (Obl., Nad.) | K | |
| 11 SkyScenes | ✓ | ✓ | ✓ | ✓ | ✓ | ✓ | ✓ | ✓ | (Low, Med, High) | (Fwd., Obl., Nad.) | K | |
1 Introduction
We introduce SkyScenes, a densely annotated dataset of synthetic scenes captured from aerial (UAV) viewpoints under diverse layout (urban and rural), weather, daytime, pitch and altitude conditions. The setting of outdoor aerial imagery provides unique challenges for scene understanding: variability in altitude and angle of image capture, skewed representation for classes with smaller object sizes (humans, vehicles), size and occlusion variations of object classes in the same image, changes in weather or daytime conditions, etc. Naturally, training effective aerial scene-understanding models requires access to large-scale annotated exemplar data that have been carefully curated under diverse conditions. Capturing such images not only allows training models that can be robust to anticipated test-time variations but also allows assessing model susceptibility to changing conditions. However, carefully curating and annotating such images in the real-world can be prohibitively expensive due to various reasons. First, densely annotating high-resolution aerial images is expensive – for instance, densely annotating a single K image in UAVid [19] can take up to hours! Second, any effort to expand the real set to include widespread variations (weather, time of day, pitch, altitude) would be uncontrolled (i.e., can’t guarantee the same viewpoint under different conditions as real world is not static) and additionally would require re-annotating newly captured frames. Synthetic data curated from simulators can help counter both of these issues as – (1) labels are automatic and cheap to obtain and (2) it is possible to recreate the same viewpoint (with the same actor instances – vehicles, humans, etc. in the scene) under differing conditions.
Unlike synthetic ground plane view datasets (especially for autonomous driving [29, 5, 37, 21, 26]), synthetic datasets for aerial imagery (see Table. 1, rows -) relatively have received less attention [9, 2, 17, 35, 16, 30, 25] Existing synthetic datasets for aerial imagery can be lacking in a few different aspects – complementary metadata to reproduce existing frame viewpoints under different conditions, limited diversity, availability of dense annotations for a wide vocabulary of classes and image capture (height, pitch) conditions (see Table. 1 for an exhaustive summary). We cover all these aspects by introducing SkyScenes, a synthetic dataset containing K densely annotated aerial scenes captured from Carla [7]. We curate SkyScenes images by re-purposing the Carla [7] simulator for aerial viewpoints by re-positioning the agent camera to a desired altitude and pitch to obtain an oblique perspective. SkyScenes images are coupled with dense semantic, instance segmentation ( classes) and depth annotations.
We carefully curate SkyScenes images by procedurally teleoperating the aerially situated camera through distinct map layouts across different weather and daytime conditions each over a combination of altitude and pitch variations (see Fig. 1 for examples). While doing so, we keep several important desiderata in mind. First, we ensure that stored snapshots are not correlated, to promote diverse viewpoints within a town and facilitate model training. Second, we store all metadata associated with the position of actors, camera, and other scene elements to be able to reproduce the same viewpoints under different weather and daytime conditions. Thirdly, we ensure that the generated data is physically realistic. This involves introducing variations in sensor locations, such as adding jitter to specified height and pitch values, to mimic real-world conditions.111Moreover, through rigorous validations, we ensure this process is consistent and yields error-free re-generations. See Sec. 3.1 Finally, since Carla [7] by default does not spawn a lot of pedestrians in a scene, we propose an algorithm to ensure adequate representation of humans in the scene while curating images (see Sec. 3.1 and Sec. 3.2 for a detailed discussion).
Empirically, we demonstrate the utility of SkyScenes in several different ways. First, we show that SkyScenes is a “good” pre-training dataset for real-world aerial scene understanding by – (1) demonstrating that models trained on SkyScenes generalize well to multiple real-world datasets and (2) that SkyScenes pretraining can help reduce data requirements from the real-world (improved real-world performance in low-shot regimes). Second, we show that controlled variations in SkyScenes can serve as a diagnostic test-bed to assess model sensitivity to weather, daytime, pitch, altitude, and layout conditions – by testing SkyScenes trained models in unseen SkyScenes conditions. Finally, we show that SkyScenes can enable developing multi-modal segmentation models with improved aerial-scene understanding capabilities when additional sensors, such as Depth, are available. To summarize, we make the following contributions:
We introduce, SkyScenes, a densely-annotated dataset of synthetic aerial images. SkyScenes contains images from different altitude and pitch settings, encompassing different layouts, weather, and daytime conditions with corresponding dense annotations and viewpoint metadata.
We demonstrate that SkyScenes pre-trained models generalize well to real-world scenes and that SkyScenes data can effectively augment real-world training data for improved performance.
We show that SkyScenes alone can serve as a diagnostic test-bed to assess model sensitivity to changing weather, daytime, pitch, altitude and layout conditions.
Finally, we show that incorporating additional modalities (depth) while training aerial scene-understanding models can improve aerial scene recognition, enabling development of multi-modal segmentation models.
2 Related Work
Ground-view Synthetic Datasets. Real world ground-view scene-understanding datasets (Cityscapes [5], Mapillary [21], BDD-100K [37], Dark Zurich [28]) fail to capture the full range of variations that exist in the world. Synthetic data is a popular alternative for generating diverse and bountiful views. GTAV [29], Synthia [26] and VisDA-C [23] are some of the widely-used synthetic datasets. These datasets can be curated using underlying simulators, such as the GTAV [29] game engine and Carla [7] simulator, and offer a cost-effective and scalable way to generate large amounts of labeled data under diverse conditions. Similar to SELMA [33] and SHIFT [32], we use Carla [7] as the underlying simulator for SkyScenes.
Real-World Aerial Datasets. To support remote sensing applications, it is crucial to have access to datasets that offer aerial-specific views. Datasets such as GID [34], DeepGlobe [6], ISPRS2D [27], and FloodNet [24] primarily provide nadir perspectives and are designed for scene-recognition and understanding tasks. However, this study specifically focuses on lower altitudes, which are more relevant to UAVs, enabling object identification. Unfortunately, there is a scarcity of high-resolution real-world datasets based on UAV imagery that emphasize on object identification. Many existing urban scene datasets, like Aeroscapes [22], UAVid [19], VDD [1], UDD [4], UAVDT [8], VisDrone [38], Semantic Drones [14] and others, suffer from limited sizes and a lack of diverse images under different conditions. This limitation raises concerns regarding model robustness and generalization.
Synthetic Aerial Datasets. Simulators can facilitate affordable, reliable, and quick collection of large synthetic aerial datasets, which aids in fast prototyping, improves real-world performance by enhancing robustness, and enables controlled studies on varied conditions. One such high-fidelity simulator, AirSim [31], used for the development and testing of autonomous systems (in particular, aerial vehicles), is the foundation of several synthetic UAV-based datasets like MidAir [9], Espada [17], Tartan Air [35], UrbanScene3d [16] and VALID [2]. Carla [7] is another such open-source simulator that is the foundation of datasets like SynDrone[25]. However, these datasets fall short in capturing real-world irregularities, lack deterministic re-generation capabilities, controlled diversity in weather and daytime conditions, and exhibit skewed representation for certain classes (differences summarized in Table. 1). This restricts their ability to generalize well to real-world datasets and their usage as a diagnostic tool for studying the controlled effect of diversity on the performance of computer vision perception tasks. To enable such studies, SkyScenes offers images featuring varied scenes, diverse weather, daytime, altitude, and pitch variations while incorporating real-world irregularities and addressing skewed class representation along with simultaneous depth, semantic, and instance segmentation annotations.
3 SkyScenes
We curate SkyScenes using, Carla [7]222https://carla.org/ , which is a flexible and realistic open-source autonomous vehicle simulator. The simulator offers a wide range of sensors, environmental configurations, and varying rendering configurations.
As noted earlier, we take several important considerations into account while curating SkyScenes images. These include strategies for obtaining diverse synthetic data and embedding real-world irregularities, avoiding correlated images, addressing skewed class representations and more. In this section, we first discuss such desiderata, and then describe our procedural image curation algorithm. Finally, we describe different aspects of the curated dataset.
3.1 (Synthetic) Aerial Image Desiderata
Before delving into the image curation pipeline, we first outline a set of desiderata taken into account while curating synthetic aerial images in SkyScenes.
1.
Adequate Height Variations: Aerial images are captured at different altitudes to meet specific needs. Lower altitudes (- m) are optimal for high-resolution photography and detailed inspections. Altitudes ranging from m-m strike a balance between fine-grained detail and a broader perspective, making them ideal for surveillance. Altitudes above m are suitable for capturing extensive areas, making them ideal for surveying and mapping. Existing datasets (synthetic or real) often focus on “specific” altitude ranges (see Table. 1, Image Capture columns), limiting their adaptability to different scenarios. With SkyScenes, our aim is to provide flexibility in altitude sampling, thus accommodating various real-world requirements. We curate SkyScenes images at heights of m, m and m. Additionally, recognizing imperfections in real-world actuation, we induce slight jitter in the height values (jitter ) to simulate realistic data sampling.
2.
Adequate Pitch Variations: Similar to height, aerial images can be captured from primary perspectives or pitch angles (): nadir (), oblique (), or forward () views (see Table. 1, Image Capture columns). The nadir view (directly perpendicular to the ground plane), preserves object scale while forward views are well-suited for tasks like UAV navigation and obstacle detection. Oblique views, on the other hand, capture objects from a side profile, aiding object recognition and providing valuable context and depth perspective often lost in nadir and forward views. To ensure widespread utility, SkyScenes data generation process is designed to support all these viewing angles, with a particular emphasis on oblique views (most common one). Similar to height, pitch variations allow models trained on SkyScenes to generalize to different real-world conditions. We use for oblique-views and introduce jitter (jitter ) to mimic real-world data sampling.
3.
Adequate Map Variations: In addition to sensor locations, it is equally important to curate aerial images across diverse scene layouts. To ensure adequate map variations, we gather images from different Carla [7] towns (can be categorized as urban or rural), which provide substantial variations in the observed scene. These towns differ in layouts, size, road map design, building design, and vegetation cover. Fig. 5 illustrates how images curated from different towns in Carla [7] differ in class distributions.
4.
Adequate Weather & Daytime Variations: Training robust perception models using SkyScenes that generalize to unforeseen environmental conditions necessitates the curation of annotated images encompassing various weather and daytime scenarios. To accomplish this, we generate SkyScenes images from identical viewpoints under different variations – ClearNoon, ClearSunset, MidRainNoon, ClearNight and CloudyNoon.333Note that Carla [7] provides such conditions but we use only such conditions in this preliminary version of SkyScenes. Generating images in different conditions from the same perspectives allows us to (1) leverage diverse data for improved generalization and (2) systematically investigate the susceptibility of trained models to variations in daytime and weather conditions.
5.
Fine-grained Annotations: To support a host of different computer vision tasks (segmentation, detection, multimodal recognition), we curate all SkyScenes images with dense semantic, instance segmentation and depth annotations. We provide semantic annotations for a wide vocabulary of classes to support broad applicability (see Fig. 1 column for an example).
Figure 2: Ground-View vs Aerial-View Pixel Proportions. For a subset of commonly annotated classes across Cityscapes [5] (red), UAVid [19] (dark blue) and SkyScenes (light blue), we show the percentage of pixels occupied by different classes. Aerial scenes (in UAVid) have significant under-representation of tail classes (vehicle, human). SkyScenes image curation (for the same viewpoint settings as UAVid) helps counter this discrepancy via synthetic aerial scenes. -
6.
Viewpoint Reproducibility: Critical to understanding how models respond to changing conditions is the ability to evaluate them under scenarios where only one variable is altered. However, any effort to do so in the real-world would be uncontrolled, due to it’s dynamic (constantly changing) nature. In contrast, simulated data allows us to do so by providing control over image generation conditions. Unlike certain existing aerial datasets that do not support this feature (see Table. 1), we do so in SkyScenes by additionally storing comprehensive metadata for each viewpoint (and image), including details about camera world coordinates, orientation and all movable / immovable actors and objects in the scene. We couple this with rigorous consistency checks for image generation that verify the number of actors, their location, sensor height, pitch, etc. This meticulous approach enables us to reproduce the same viewpoint under multiple conditions effortlessly.
7 .
Adequate Representation of Tail Classes: Unlike ground-view datasets, pixel distribution of classes in aerial images is substantially more long-tailed (see Fig. 2; classes with smaller object size, humans), making visual recognition tasks harder. This problem is further exacerbated by differing class distributions across synthetic and real data (see Fig. 5) – especially severe for instances of human class. To counter this, we consider structured spawning of humans to ensure adequate representation.
3.2 SkyScenes Image Generation
We generate SkyScenes images from Carla [7] by taking the previously mentioned considerations into account. Curating images from Carla [7] broadly consists of two key steps: (1) positioning the agent camera in an aerial perspective and (2) procedurally guiding the agent within the scene to capture images. We accomplish the first by mimicking a UAV perspective in Carla [7] by positioning the ego vehicle (with RGB, semantic and depth sensors) based on specified (high) altitude () and pitch () values to generate aerial views (see Fig. 3).444This also requires setting other scene – weather, daytime, etc. – and camera – notably the (field of view) and (image resolution) – parameters. Once positioned, the agent is translated by fixed amounts to traverse the scene and capture images from various viewpoints (detailed in Algo. 1 in the appendix). Initially, we generate datapoints for each of the town variations under ClearNoon conditions using the baseline setting. Subsequently, following
the traversal algorithm (see Algo. 1 in the appendix),
we re-generate these datapoints across weather conditions and height/pitch variations,
resulting in images.
Checks and Balances.
Additionally, we ensure the following checks and balances while curating SkyScenes images.
Avoiding Overly Correlated Frames for Viewpoints.
Carla [7] uses a traffic manager with a PID controller to control the egocentric vehicle based on current pose, speed, and a list of waypoints at every pre-defined time step. Curating images at every time step (or tick) results in highly correlated frames with little change in object position. Since overly correlated frames are not very useful when training models for static scene understanding, we move the camera by a fixed distance multiple times before saving a frame. This also helps with moving dynamic actors by a considerable amount in the scene. Additionally, pedestrian objects are regenerated before saving an image, which adds randomness to the spawning and placement of pedestrians and further reduces correlation between frames.
| Layout | RS (%) | HS (%) |
| 1 Town01 | ||
| 2 Town02 | ||
| 3 Town03 | ||
| 4 Town04 | ||
| 5 Town05 | ||
| 6 Town06 | ||
| 7 Town07 | ||
| 8 Town10HD | ||
| 9 All | ||
| 10 UAVid [19] |
| mIoU | |||
| Eval Data | HS | human | All |
| 1 S | |||
| 2 S | ✓ | 61.79 | 84.07 |
| 3 SU | |||
| 4 SU | ✓ | 10.21 | 47.09 |
Adequate Representation of humans. Real-world scenes often exhibit a long-tailed distribution in pixel proportions, particularly in aerial images where variations in object sizes and camera positions contribute to significant under-representation of the tail classes (in Fig. 2, for the shared set of classes across UAVid [19] (aerial) and Cityscapes [5] (ground), we can see that the class distributions are different and aerial images are significantly more heavy tailed). As a result, naively spawning humans (rarest class) in Carla [7] is detrimental for eventual task performance – for the human class, a SkyScenes trained DAFormer [11](with HRDA [12] source training; MiT-B5 [36] backbone) model leads to an in-distribution performance of mIoU and out-of-distribution (SkyScenes UAVid [19]) performance of mIoU. To counter this under-representation issue, we design an algorithm, HumanSpawn()555More details in Sec. B.1 of the appendix., to explicitly spawn more human instances while curating SkyScenes images. HumanSpawn() increases human instances by per snapshot, improving the proportion of densely annotated humans in SkyScenes by approximately times (see Table. 3 & Fig. 4). This improvement in human representation is also evident in eventual task performance, with in-distribution and out-of-distribution mIoUs for humans increasing from to () and to () respectively (see Table. 3.)
3.3 SkyScenes: Dataset Details
Annotations. We provide semantic, instance and depth annotations for every image in SkyScenes. Semantic annotations in SkyScenes by default are across classes. These are building, fence, pedestrian, pole, roadline (markings on road), road, sidewalk, vegetation, cars, wall, traffic sign, sky, bridge, railtrack, guardrail, traffic light, water, terrain, rider, bicycle, motorcycle, bus, truck and others (see Fig. 1 for an example).666We provide detailed definitions in Sec. B.2 of the appendix.
Training, Validation and Test Splits. SkyScenes has images per town (across towns) for each of the weather and daytime conditions, and height & pitch combinations, resulting in a total of images. We use % ( images) of the dataset for training models, with % ( images) each for validation and testing. While creating train, val and test splits, we collect equal number of samples from each town by dividing each town-specific traversal sequence into segments: the initial % for training, the next % for testing, and the final % of the segment for validation. Moreover, within each split, we ensure that every viewpoint is accompanied by its different variations across weather, daytime, height, and pitch settings. This safeguards against any potential cross-contamination across different splits while ensuring fair representation and equal distributions of all variations.
Class Distribution(s). In Fig. 6, we compare the number of densely annotated pixels in SkyScenes with those in Syndrone [25] (another synthetic dataset) and UAVid [19] (a real aerial dataset). Compared to both the real and prior synthetic counterparts, we show that SkyScenes is specifically curated to ensure adequate tail class representation (more human and vehicle pixels) to facilitate model learning. Additionally, in Fig. 5, we highlight how the distribution of classes changes across variations within SkyScenes– rural and urban map layouts and height and pitch specifications. SkyScenes exhibits substantial diversity in class-distributions across such conditions, allowing these individual conditions to serve as diagnostic splits to assess model sensitivity (see Sec. 4.2)
4 Experiments
We conduct semantic segmentation experiments with SkyScenes to assess a few different factors. First, we check if training on SkyScenes is beneficial for real-world transfer. Second, we check if SkyScenes can augment real-world training data in low and full shot regimes. Third, we check if variations in SkyScenes can be used to assess sensitivity of trained models to changing conditions. Finally, we check if using additional modality information (depth) can help improve aerial scene understanding.
Synthetic and Real Datasets. We compare real-world generalization performance of training on SkyScenes with SynDrone [25], a recently proposed synthetic aerial dataset also curated from Carla [7]. We assess performance on real-world aerial datasets – UAVid [19], AeroScapes [22], ICG Drone [14]. Since different datasets have different class vocabularies and definitions, for our experiments, we adapt the class vocabulary of the synthetic source dataset to that of the target real-world datasets.777We detail class merging and assignment schemes used for these experiments in Table. 8, Table. 9 and Table. 10 in the appendix. Additionally, since different real aerial datasets have been captured from different heights and pitch angles, we train models on subsets of synthetic datasets that are aligned with corresponding real data conditions. We provide additional details for the real aligned synthetic data selection and model evaluation in Sec. D.1 of the appendix.
Models. We use both CNN – DeepLabv2 [3] (ResNet-101 [10]) – and transformer – DAFormer [11] (with HRDA [12] source training; MiT-B5 [36] backbone) – based semantic segmentation architectures for our experiments. We provide implementation details surrounding our experiments in Sec. C of the appendix.
| Source | (Target) Real-World mIoU () | ||
| UAVid | AeroScapes | ICG Drone | |
| DeepLabv2 (R-101) [3] | |||
| 1 SynDrone | |||
| 2 SkyScenes | 41.82 | 26.94 | 15.14 |
| 3 Train-on-Target | |||
| DAFormer (MiT-B5) [12] | |||
| 4 SynDrone | |||
| 5 SkyScenes | 47.09 | 40.72 | 25.91 |
| 6 Train-on-Target | |||
| Source | (Target) Real-World mIoU () | |||||
| UAVid | AeroScapes | ICG Drone | ||||
| vehicle | human | vehicle | person | vehicle | person | |
| 1 SynDrone | ||||||
| 2 SkyScenes | 63.64 | 10.21 | 80.99 | 3.09 | 39.71 | 45.89 |
| 3 Train-on-Target | ||||||
4.1 SkyScenes Real Transfer
SkyScenes trained models generalize well to real-settings. In Table. 4, we show how models trained on SkyScenes exhibit strong out-of-the box generalization performance on multiple real world datasets. We find that SkyScenes pretraining exhibits stronger generalization compared to SynDrone [25] across both CNN and transformer segmentation backbones. In Table. 5, we show how generalization improvements are more pronounced for under-represented tail classes (vehicles and humans)888Model performance comparisons for all the classes are provided in Table 11, Table. 12 and Table. 13 in the appendix.
SkyScenes can augment real training data. In addition to zero-shot real-world generalization, akin to other synthetic aerial datasets, we also show how SkyScenes is useful as additional training data when labeled real-world data is available. In Fig. 7, for SkyScenes UAVid [19], we compare models trained only using of the UAVid [19] training images with counterparts that were either pretrained using SkyScenes data or additionally supplemented with SkyScenes data at training time. We find that in low-shot regimes (when little “real” world data is available), SkyScenes data (either explicitly via joint training or implicitly via finetuning) is beneficial in improving recognition performance. We find this to be especially beneficial for under-represented classes in aerial imagery (such as humans and vehicles). We discuss more such results across other real-world datasets in Sec. D.2 of the appendix.
| Train | Test mIoU () | ||
| Clear | Cloudy | Rainy | |
| 1 Clear | |||
| 2 Cloudy | |||
| 3 Rainy | |||
| Train | Test mIoU () | ||
| Noon | Sunset | Night | |
| 1 Noon | |||
| 2 Sunset | |||
| 3 Night | |||
| Train | Test mIoU () | |
| Rural | Urban | |
| 1 Rural | ||
| 2 Urban | ||
| Test mIoU () | ||||
| Height | Pitch | |||
| 1 m | 48.50 | |||
| 2 m | ||||
| 3 m | ||||
4.2 SkyScenes as a Diagnostic Framework
As noted earlier, the images we curate in SkyScenes contain several variations – ranging from different weather and daytime conditions, rural and urban map layouts, and different height and pitch combinations (see Fig. 5 for variations in class distributions). We curate images under such diverse conditions in a controlled manner – ensuring the same spatial coordinates for variations, same spatial coordinates and settings across different weather and daytime conditions, same number of images across layouts. This allows us to assess the sensitivity of trained models to one factor of variation (, , daytime, weather, map layout) by changing that specific aspect. We summarize some takeaways from such experiments in Table. 6.
In Table. 6 (a), we show how models trained in a certain weather condition are best at generalizing to the same condition at test-time. We make similar observations for daytime variations in Table. 6 (b). In Table. 6 (c), we show how models trained in rural conditions fail to perform well in urban test-time conditions and vice-versa. In Table. 6 (d), we evaluate a model trained under moderate conditions under different variations. We find that as altitudes increase, trained models are better at recognizing objects from oblique () viewpoints. We provide exhaustive quantitative comparisons in Sec. D.3 of the appendix.
| Sensors | SkyScenes Test mIoU () | |||||||
| clutter | building | road | tree | low-veg. | vehicle | human | Avg | |
| 1 RGB | ||||||||
| 2 RGB+D | 90.64 | 95.97 | 94.87 | 89.41 | 74.36 | 86.87 | 50.47 | 83.22 |
4.3 SkyScenes Enables Multi-modal Recognition
Sensors on UAVs in deployable settings are not limited to RGB cameras. It is common to have UAVs deployed in the real-world with additional modality sensors, such as depth. Additional sensor modalities can also potentially help improve aerial scene understanding. In Table. 7, we check if assisting RGB with Depth observations for SkyScenes viewpoints can help improve aerial semantic segmentation using M3L [20], a model capable of multimodal segmentation. Similar to our DAFormer [11] experiments, we consider a SegFormer equivalent version of M3L [20] (with an MiT-B5 [36] backbone). We test RGB and RGB+D models trained under (moderate viewpoint) conditions on SkyScenes and find that incorporating additional Depth observations can substantially improve recognition performance. This demonstrates that annotated images in SkyScenes can be used to train multimodal scene-recognition models.
5 Conclusion
We introduce SkyScenes, a large-scale densely-annotated dataset of synthetic aerial scenes curated from unmanned aerial vehicle (UAV) perspectives. We collect SkyScenes images from Carla by first aerially situating an agent (to get an aerial perspective) and then procedurally tele-operating the agent through the scene to capture aerial frames with corresponding semantic, instance and depth annotations. Our careful curation process ensures that SkyScenes images are carefully curated across diverse weather, daytime, map, height and pitch conditions, with accompanying metadata that enables reproducing the same viewpoint (spatial coordinates and perspective) under differing conditions.
Through our experiments, we demonstrate how (1) SkyScenes trained models can generalize to real-world settings, (2) SkyScenes images can augment labeled real-world data in low-shot regimes, (3) SkyScenes can serve as a diagnostic framework to assess model sensitivity to changing conditions and (4) additional sensors, such as Depth, in SkyScenes can facilitate development of multi-modal aerial scene understanding models.
Lastly, we plan on updating SkyScenes with evolving considerations for real-world aerial scene-understanding – improved realism, additional anticipated edge cases – as more and more features are supported in the underlying simulator. We intend to release both the dataset and associated generation code for SkyScenes publicly, and hope that our experimental findings encourage further research using SkyScenes for aerial scenes.
Acknowledgements. We would like to thank Sean Foley for his contributions to the early efforts and discussions of this project. This work has been partially sponsored by NASA University Leadership Initiative (ULI) #80NSSC20M0161, ARL, and NSF #2144194.
Appendix A Appendix
This appendix is organized as follows. In Sec. B, we provide more details on different aspects of dataset – including the procedural image curation algorithm, algorithm to ensure appropriate representation of human instances and brief descriptions of all the classes in SkyScenes. Then, we describe experimental details in Sec. C – class-merging schemes used for our SkyScenes Real transfer experiments, train / val / test splits, experimental details for SkyScenes diagnostic setup to probe model vulnerabilities. Sec. D provides more quantitative and qualitative experimental results.
Appendix B SkyScenes Details
B.1 Image Generation Algorithms
We curate SkyScenes using Carla [7] simulator. The generation process broadly consists of two key steps: (1) positioning the agent camera in an aerial perspective and (2) procedurally guiding the agent within the scene to capture images. We accomplish the first by mimicking a UAV perspective in Carla [7] by positioning the ego vehicle (with RGB, semantic, depth and instance segmentation sensors) based on specified altitude () and pitch () values to generate aerial views. This also requires setting other scene information like – town, weather, and daytime. The camera (field of view) and (image resolution) are also set.
To maintain the adequate representation of tail classes, especially for humans we incorporate structured spawning of humans using Algo. 2. Carla [7] has a limit on the number of actors that can be spawned in a scene, which depends on factors such as the type and size of the town, number of lanes to spawn vehicles, and amount of sidewalk area. To overcome this limitation, we decided to bring the actors into the field of view of the camera instead of having them spread out in the scene. We developed an algorithm to find all the points to spawn pedestrians in the field of view using the camera location and spawn them like vehicles on roads. After taking a snapshot, we destroy the spawned pedestrians and repeat the process. Manual spawning not only increases the number of human instances and their proportion but also aligns their placement with real-world settings. The steps involved in manual spawning instances of humans are summarized below:
1
Specify maximum number of humans to be spawned
2
Get camera position , set a pre-defined distance to check for spawnable locations and execute the subroutine in Algo. 2. This will place the actors in the field of view till a junction or the next driving lane.
3
If at a junction, obtain the left and the right waypoints for every retrieved location at a distance of to get the list of waypoints.
4
Using the waypoint from the end of the current lane, generate waypoints for the new main, left and right lanes by repeating the previous steps.
5
Repeat the above steps till
Once positioned, the agent is translated by fixed amounts to traverse the scene and capture images from various viewpoints. Initially, we generate datapoints for each of the town variations under ClearNoon conditions using the baseline setting. Subsequently, following the traversal algorithm (Algo. 1), we re-generate these datapoints across weather conditions and height/pitch variations, resulting in images.
B.2 Class Descriptions
We provide semantic, instance, and depth annotations for every image in SkyScenes. SkyScenes provides dense semantic annotations for classes. These are:
unlabeled: elements/objects in the scene that have not been categorized in Carla
other: uncategorized elements
building: includes houses, skyscrapers, and the elements attached to them.
fence: wood or wire assemblies that enclose an area of ground
pedestrian: humans that walk
pole: vertically oriented pole and its horizontal components if any
roadline: markings on road.
road: lanes, streets, paved areas on which cars drive
sidewalk: parts of ground designated for pedestrians or cyclists
vegetation: trees, hedges, all kinds of vertical vegetation (ground-level vegetation is not included here).
cars: cars in scene
wall: individual standing walls, not part of buildings
traffic sign: signs installed by the state/city authority, usually for traffic regulation
sky: open sky, including clouds and sun
ground: any horizontal ground-level structures that does not match any other category
bridge: the structure of the bridge
railtrack: rail tracks that are non-drivable by cars
guardrail: guard rails / crash barriers
traffic light: traffic light boxes without their poles.
static: elements in the scene and props that are immovable.
dynamic: elements whose position is susceptible to change over time.
water: horizontal water surfaces
terrain: grass, ground-level vegetation, soil or sand
rider: humans that ride/drive any kind of vehicle or mobility system
bicycle: bicycles in scenes
motorcycle: motorcycles in scene
bus: buses in scenes
truck: trucks in scenes
B.3 Depth & Instance Segmentation
We also provide depth and instance segmentation annotations along with semantic segmentation annotations. Both the depth and instance segmentation sensors are mounted alongside the RGB camera and semantic segmentation sensors. Depth is stored in the format which provides better results for closer objects. We also provide depth-aided semantic segmentation results (training details in Sec. C.3).
Appendix C Experiment Details
C.1 Class Merging Details
| Syn UAVid | UAVid | SkyScenes | SynDrone |
| 1 clutter | clutter | unlabelled | unlabelled |
| other | other | ||
| traffic sign | traffic sign | ||
| rail track | rail track | ||
| guard rail | guard rail | ||
| traffic light | traffic light | ||
| static | static | ||
| dynamic | dynamic | ||
| ground | ground | ||
| sidewalk | sidewalk | ||
| fence | fence | ||
| sky | sky | ||
| bridge | bridge | ||
| water | water | ||
| pole | pole | ||
| 2 building | building | building | building |
| wall | wall | ||
| 3 road | road | road | road |
| roadline | roadline | ||
| 4 vegetation | vegetation | vegetation | vegetation |
| 5 low vegetation | terrain | terrain | terrain |
| 6 person | person | pedestrian | pedestrian |
| rider | rider | ||
| motorcycle | motorcycle | ||
| bicycle | bicycle | ||
| 7 vehicle | car | car | car |
| truck | truck | ||
| bus | bus | ||
| train |
| Syn AeroScapes | AeroScapes | SkyScenes | SynDrone |
| 1 background | background | unlabelled | unlabelled |
| drone | other | other | |
| boat | traffic sign | traffic sign | |
| animal | rail track | rail track | |
| obstacle | guard rail | guard rail | |
| traffic light | traffic light | ||
| static | static | ||
| dynamic | dynamic | ||
| ground | ground | ||
| sidewalk | sidewalk | ||
| terrain | terrain | ||
| water | water | ||
| pole | pole | ||
| 2 bicycle | bike | bicycle | bicycle |
| motorcycle | motorcycle | ||
| 3 person | person | pedestrian | pedestrian |
| rider | rider | ||
| 4 vehicle | car | car | car |
| truck | truck | ||
| bus | bus | ||
| train | |||
| 5 vegetation | vegetation | vegetation | vegetation |
| 6 building | construction | building | building |
| wall | wall | ||
| fence | fence | ||
| bridge | bridge | ||
| 7 road | road | road | road |
| roadline | roadline | ||
| 8 sky | sky | sky | sky |
| Syn ICG Drone | ICG Drone | SkyScenes | SynDrone |
| 1 other | obstacle | unlabelled | unlabelled |
| dog | other | other | |
| conflicting | traffic sign | traffic sign | |
| ar-marker | sky | sky | |
| unlabelled | bridge | bridge | |
| gurad rail | guard rail | ||
| traffic light | traffic light | ||
| static | static | ||
| dynamic | dynamic | ||
| 2 fence | fence | fence | fence |
| 3 pole | fence pole | pole | pole |
| 4 vegetation | tree | vegetation | vegetation |
| bald-tree | |||
| 5 building | wall | wall | wall |
| roof | building | building | |
| door | |||
| window | |||
| 6 water | water | water | water |
| pool | |||
| 7 bicycle | bicycle | bicycle | bicycle |
| motorcycle | motorcycle | ||
| 8 vehicle | car | car | car |
| truck | truck | ||
| bus | bus | ||
| train | |||
| 9 person | person | pedestrian | pedestrian |
| rider | rider | ||
| 10 paved area | paved area | ground | ground |
| sidewalk | sidewalk | ||
| road | road | ||
| roadline | roadline | ||
| rail track | rail track | ||
| 11 terrain | rocks | terrain | terrain |
| gravel | |||
| dirt | |||
| vegetation | |||
| grass |
SyntheticReal. As noted in Sec of the main paper, since different real-world datasets have different class vocabularies and definitions, for our SyntheticReal semantic segmentation experiments, we adapt the class-vocabulary of the synthetic source dataset (SkyScenes, SynDrone) to that of the target real dataset (UAVid, AeroScapes, ICG Drone). This is done using a class-merging scheme based on the class-vocabularies and after visually inspecting dataset annotations. We provide the class-merging schemes used for both SkyScenes and SynDrone [25] (synthetic) across real counterparts UAVid [19] (in Table. 8), AeroScapes [22] (in Table. 9) and ICG Drone [14] (in Table. 10).
SkyScenes Diagnostic Experiments. To assess the sensitivity of trained models to different factors – weather, time of day, height, pitch, etc.– we train models on different SkyScenes variations and evaluate them on held-out-variations. For these experiments, we recude the SkyScenes vocabulary to a reasonable subset of classes (consistent with the widely used Cityscapes [5] palette) – road, sidewalk, building, wall, fence, pole, traffic light, traffic sign, vegetation, water, sky, pedestrian, rider, cars, truck, bus, roadline, motorcycle, bicycle and an ignore class.
C.2 Training, Validation and Test Splits
C.2.1 Synthetic Real Experiments
SkyScenes. For each
combination,
SkyScenes has a total of datapoints (frames) which are distributed evenly across each of the town layouts and weather and daytime conditions. We use % ( images) of these datapoints for training models,
and remaining
% ( images) each for validation and testing.
SynDrone. SynDrone has images per town (across Carla towns) for each of the
combinations,
resulting in a total of images per
combination.
We use of these datapoints for training models, with kept aside for testing. The datapoints selected for training and testing are kept consistent with the one reported in SynDrone [25].
C.2.2 SkyScenes Diagnostic Experiments
Weather & Daytime Variations. For each weather variation we sample evenly across the towns and
combinations (excluding ),
resulting in a total of images.
We use
% ( images) of these
for training models, and remaining % ( images) each for validation and testing. We evaluate the model on the same weather variation it was trained on and select the model with the best mIoU score for further evaluations on other weather variations.
Town Variations. For each variation in town we sample evenly across weather and daytime conditions and height and pitch variations(excluding pitch= variations), resulting in a total of images per town variation. Out of these, % ( images) is allocated for training models, with % ( images) each for validation and testing. We evaluate the model on the same town variation it was trained on and select the model with the best mIoU score for further evaluations on other town variations.
Height & Pitch Variations. For each variation, we evenly sample across towns and weather and daytime conditions, resulting in a total of images per height and pitch variation. Out of these, % ( images) are allocated for training models, and % ( images) each for validation and testing.We evaluate the model on the same height & pitch setting it was trained on and select the model with the best mIoU score for further evaluations on other height & pitch settings.
C.3 Training Details
Semantic Segmentation. For our semantic segmentation experiments we use both CNN – DeepLabv2 [3] (ResNet-101 [10] backbone) – and vision transformer – DAFormer [11] (with HRDA [12] source training; MiT-B5 [36] backbone) – based semantic segmentation architectures. Note that we utilize only the DAFormer architecture to perform source-only training for our experiments. Following [13], we enable rare class sampling [11] and use Imagenet feature-distance for our thing classes during training. All models are trained using the AdamW [18] optimizer coupled with a polynomial learning rate scheduler with an initial learning rate of . Each model is trained for k iterations with a batch size of . For our fine-tuning experiments, we use an initial learning rate of .
Depth-Aided Semantic Segmentation. For our depth-aided semantic segmentation experiments in Sec. 4.3 of the main paper, similar to DAFormer [11], we employ a SegFormer [36] equivalent version of M3L [20] (multimodal segmentation network) with an MiT-B5 [36] backbone. We initialize the network with ImageNet-1k pre-trained checkpoints. For M3L Linear Fusion, we use . We use AdamW [15] optimizer and train on a batch size of for epochs. We use a learning rate of for the encoder and for the decoder with a momentum of and weight decay of . We set the polynomial decay of power . We train both RGB and RGB+D models with complete supervision for (moderate viewpoint) conditions on SkyScenes.
C.4 Evaluation Details
Due to memory constraints, in addition to heavily parameterized models, our GPUs were unable to fit images larger than . Hence for high-resolution datasets like UAVid, we use the trained model to make separate predictions on equally sized slightly-overlapping crops (overlap of pixels) of the of the real image and stitch crop predictions to obtain the overall image prediction. Similarly, for ICG Drone, we obtain overall image prediction using such crop predictions.
Appendix D Results
D.1 SyntheticReal Aligned Data Selection
As stated in the main paper, for our SyntheticReal experiments, we train models on subsets of synthetic datasets that are aligned with corresponding real data conditions. In case of UAVid and AeroScapes we find that viewpoints in SkyScenes and viewpoints in SynDrone are best aligned with UAVid conditions (see Table. 11 and Table. 12) and provide best transfer performance. Similarly, for ICG Drone we observe that SkyScenes conditions are best aligned with the low-altitude, nadir perspective imagery in ICG Drone and lead to best transfer performance (see Table. 13). However, for SynDrone, we find that the model trained on has best transfer performance, indicating that model performance is more sensitive to pitch alignment than height alignment.
| h(m) | Synthetic mIoU () | ||||||||
| Clutter | Building | Road | Tree | Low Vegetation | Human | Vehicle | Avg | ||
| SkyScenes | |||||||||
| 1 | |||||||||
| 2 | |||||||||
| 3 | |||||||||
| 4 | |||||||||
| 5 | |||||||||
| 6 | |||||||||
| 7 | |||||||||
| 8 | |||||||||
| 9 | |||||||||
| 10 | |||||||||
| 11 | |||||||||
| 12 | |||||||||
| SynDrone | |||||||||
| 13 | |||||||||
| 14 | |||||||||
| 15 | |||||||||
| h(m) | Synthetic mIoU () | |||||||||
| Background | Bicycle | Person | Vehicle | Vegetation | Building | Road | Sky | Avg | ||
| SkyScenes | ||||||||||
| 1 | ||||||||||
| 2 | ||||||||||
| 3 | ||||||||||
| 4 | ||||||||||
| 5 | ||||||||||
| 6 | ||||||||||
| 7 | ||||||||||
| 8 | ||||||||||
| 9 | ||||||||||
| 10 | ||||||||||
| 11 | ||||||||||
| 12 | ||||||||||
| SynDrone | ||||||||||
| 13 | ||||||||||
| 14 | ||||||||||
| 15 | ||||||||||
| h(m) | Synthetic mIoU () | |||||||||||||
| other | fence | pole | vegetation | building | water | bicycle | vehicle | person | paved area | terrain | Avg | Avg | ||
| SkyScenes | ||||||||||||||
| 1 | ||||||||||||||
| 2 | ||||||||||||||
| 3 | ||||||||||||||
| 4 | ||||||||||||||
| 5 | ||||||||||||||
| 6 | ||||||||||||||
| 7 | ||||||||||||||
| 8 | ||||||||||||||
| 9 | ||||||||||||||
| 10 | ||||||||||||||
| 11 | ||||||||||||||
| 12 | ||||||||||||||
| SynDrone | ||||||||||||||
| 13 | ||||||||||||||
| 14 | ||||||||||||||
| 15 | ||||||||||||||
| Training | Data Size | Synthetic mIoU () | |||||||
| Clutter | Building | Road | Tree | Low Vegetation | Human | Vehicle | Avg | ||
| SkyScenes | |||||||||
| 1 FT | 53.67 | ||||||||
| 2 FT | 60.61 | ||||||||
| 3 FT | 65.57 | ||||||||
| 4 FT | 68.54 | ||||||||
| 5 FT | 69.70 | ||||||||
| 6 JT | 59.27 | ||||||||
| 7 JT | 64.15 | ||||||||
| 8 JT | 68.11 | ||||||||
| 9 JT | 70.18 | ||||||||
| 10 JT | |||||||||
| SynDrone | |||||||||
| 11 FT | |||||||||
| 12 FT | |||||||||
| 13 FT | |||||||||
| 14 FT | |||||||||
| 15 FT | |||||||||
| 16 JT | |||||||||
| 17 JT | |||||||||
| 18 JT | |||||||||
| 19 JT | |||||||||
| 20 JT | 70.01 | ||||||||
| Target | |||||||||
| 21 Target | |||||||||
| 22 Target | |||||||||
| 23 Target | |||||||||
| 24 Target | |||||||||
| 25 Target | |||||||||
| Training | Data Size | Synthetic mIoU () | |||||||
| Clutter | Building | Road | Tree | Low Vegetation | Human | Vehicle | Avg | ||
| SkyScenes | |||||||||
| 1 FT | |||||||||
| 2 FT | |||||||||
| 3 FT | |||||||||
| 4 FT | |||||||||
| 5 FT | |||||||||
| 6 JT | 62.97 | ||||||||
| 7 JT | 67.58 | ||||||||
| 8 JT | 70.20 | ||||||||
| 9 JT | 71.83 | ||||||||
| 10 JT | |||||||||
| SynDrone | |||||||||
| 11 FT | |||||||||
| 12 FT | |||||||||
| 13 FT | |||||||||
| 14 FT | |||||||||
| 15 FT | |||||||||
| 16 JT | |||||||||
| 17 JT | |||||||||
| 18 JT | |||||||||
| 19 JT | |||||||||
| 20 JT | |||||||||
| Target | |||||||||
| 21 Target | |||||||||
| 22 Target | |||||||||
| 23 Target | |||||||||
| 24 Target | |||||||||
| 25 Target | |||||||||
| Training | Data Size | Synthetic mIoU () | ||||||||
| Background | Bicycle | Person | Vehicle | Vegetation | Building | Road | Sky | Avg | ||
| SkyScenes | ||||||||||
| 1 FT | 62.57 | |||||||||
| 2 FT | 68.28 | |||||||||
| 3 FT | 69.84 | |||||||||
| 4 FT | 72.47 | |||||||||
| 5 FT | 72.71 | |||||||||
| 6 JT | 69.16 | |||||||||
| 7 JT | 74.10 | |||||||||
| 8 JT | 74.91 | |||||||||
| 9 JT | ||||||||||
| 10 JT | 77.13 | |||||||||
| SynDrone | ||||||||||
| 11 FT | ||||||||||
| 12 FT | ||||||||||
| 13 FT | ||||||||||
| 14 FT | ||||||||||
| 15 FT | ||||||||||
| 16 JT | ||||||||||
| 17 JT | ||||||||||
| 18 JT | ||||||||||
| 19 JT | 76.22 | |||||||||
| 20 JT | ||||||||||
| Target | ||||||||||
| 21 Target | ||||||||||
| 22 Target | ||||||||||
| 23 Target | ||||||||||
| 24 Target | ||||||||||
| 25 Target | ||||||||||
| Training | Data Size | Synthetic mIoU () | ||||||||
| Background | Bicycle | Person | Vehicle | Vegetation | Building | Road | Sky | Avg | ||
| SkyScenes | ||||||||||
| 1 FT | ||||||||||
| 2 FT | ||||||||||
| 3 FT | 76.56 | |||||||||
| 4 FT | 77.93 | |||||||||
| 5 FT | 78.31 | |||||||||
| 6 JT | 71.94 | |||||||||
| 7 JT | 76.53 | |||||||||
| 8 JT | ||||||||||
| 9 JT | ||||||||||
| 10 JT | 80.06 | |||||||||
| SynDrone | ||||||||||
| 11 FT | ||||||||||
| 12 FT | ||||||||||
| 13 FT | ||||||||||
| 14 FT | ||||||||||
| 15 FT | ||||||||||
| 16 JT | ||||||||||
| 17 JT | ||||||||||
| 18 JT | ||||||||||
| 19 JT | ||||||||||
| 20 JT | ||||||||||
| Target | ||||||||||
| 21 Target | ||||||||||
| 22 Target | ||||||||||
| 23 Target | ||||||||||
| 24 Target | ||||||||||
| 25 Target | ||||||||||
D.2 SkyScenes Real Data Experiments
In addition to zero-shot transfer to real data, we also show how SkyScenes is useful as additional training data when labeled real-world data is available. In Fig. 8 and Fig. 9, we compare the performance of DeepLabv2 [3] for SkyScenes UAVid [19] and for SkyScenes AeroScapes [22] trained only using of UAVid [19] and AeroScapes [22] training images respectively with counterparts that were either pretrained using SkyScenes data or additionally supplemented with SkyScenes data at training time. In Fig. 10 and Fig. 11 we make a similar comparison with the DAFormer [11] architecture. In low-shot regimes (when little “real” world data is available), SkyScenes data (either explicitly via joint training or implicitly via finetuning) is beneficial in improving recognition performance. We find this to be especially beneficial for under-represented classes in aerial imagery (such as humans and vehicles). In Table. 13(a) and Table. 13(c) we present similar fine-grained (per-class) comparison of SkyScenes with SynDrone for a DeepLabv2 model when real data is available for training – via Target-Only, Finetuning or Joint-Training. Tables. 13(b) and 13(d) show similar comparisons using the DAFormer [11] model.
D.3 SkyScenes Diagnostic Experiments
Similar to Table 6 (d) in the main paper, in Table 15, we assess broader sensitivity of models by training DAFormer [11] models across all (total ) settings and evaluate them across the same conditions in SkyScenes. Table 15 (a) models trained on are representative of one extreme of the range (both lowest height and pitch values) – we notice that this model is extremely sensitive to height variations. From Tables 15 (a), (e) and (i), we can deduce that due to the high variability in perspective between and other conditions, models trained on do not generalize well to other values. On the other hand, Tables. 15 (b), (c), (d), (f), (g) and (h) models trained on oblique perspectives are better at generalizing to other pitch conditions.
| Test mIoU () | ||||
| Height | Pitch | |||
| 1 m | 72.57 | 66.38 | 57 | 39.46 |
| 2 m | 59.07 | 55.55 | 54.43 | 39.23 |
| 3 m | 48.58 | 44.94 | 43.85 | 31.71 |
| Test mIoU () | ||||
| Height | Pitch | |||
| 1 m | 59.03 | 66.98 | 63.07 | 61.85 |
| 2 m | 43.8 | 49.41 | 49.91 | 45.19 |
| 3 m | 34.27 | 39.86 | 40.1 | 35.42 |
| Test mIoU () | ||||
| Height | Pitch | |||
| 1 m | 45.36 | 64.13 | 68.88 | 69.28 |
| 2 m | 30.49 | 45.05 | 50.39 | 53.44 |
| 3 m | 21.86 | 31.21 | 36.51 | 38.43 |
| Test mIoU () | ||||
| Height | Pitch | |||
| 1 m | 30.61 | 55.79 | 64.98 | 71.93 |
| 2 m | 16.84 | 33.2 | 37.89 | 48.63 |
| 3 m | 11.26 | 21.12 | 26.38 | 31.22 |
| Test mIoU () | ||||
| Height | Pitch | |||
| 1 m | 52.14 | 47.62 | 39.04 | 29.31 |
| 2 m | 58.03 | 55.61 | 55.07 | 45.54 |
| 3 m | 53.31 | 50.38 | 50.23 | 46.65 |
| Test mIoU () | ||||
| Height | Pitch | |||
| 1 m | 48.50 | |||
| 2 m | ||||
| 3 m | ||||
| Test mIoU () | ||||
| Height | Pitch | |||
| 1 m | 37.21 | 49.04 | 47.42 | 44.93 |
| 2 m | 34.37 | 52.67 | 57.52 | 54.14 |
| 3 m | 29.82 | 44.36 | 48.33 | 44.71 |
| Test mIoU () | ||||
| Height | Pitch | |||
| 1 m | 28.73 | 40.11 | 44.9 | 46 |
| 2 m | 25.89 | 38.98 | 46.26 | 54.02 |
| 3 m | 20.36 | 29.95 | 36.1 | 43.16 |
| Test mIoU () | ||||
| Height | Pitch | |||
| 1 m | 37.59 | 32.89 | 24.53 | 18.35 |
| 2 m | 48.42 | 44.31 | 42.82 | 32.04 |
| 3 m | 51.53 | 48.38 | 47.71 | 39.45 |
| Test mIoU () | ||||
| Height | Pitch | |||
| 1 m | 43.01 | 43.33 | 35.6 | 31.05 |
| 2 m | 52.13 | 57.6 | 58.84 | 51.95 |
| 3 m | 51.89 | 56.83 | 56.43 | 50.07 |
| Test mIoU () | ||||
| Height | Pitch | |||
| 1 m | 38.58 | 43.4 | 36.96 | 32.22 |
| 2 m | 44.69 | 57.52 | 60.21 | 54.3 |
| 3 m | 43.24 | 56.68 | 59.27 | 49.18 |
| Test mIoU () | ||||
| Height | Pitch | |||
| 1 m | 28.12 | 34.9 | 33.6 | 33.28 |
| 2 m | 32.76 | 45.32 | 53.71 | 55.27 |
| 3 m | 29.58 | 42.59 | 50.73 | 57.07 |
D.4 Qualitative examples
In Fig. 12, 13 and 14, we show qualitative examples of predictions made by SkyScenes and SynDrone trained DeepLabv2 and DAFormer models on UAVid, AeroScapes, and ICG Drone respectively. In Fig. 15 and 16, we show how predictions are impacted by changing SkyScenes conditions.

### v2

@@
 Due to inherent challenges in obtaining such images in controlled real-world settings, we present SkyScenes, a synthetic dataset of densely annotated aerial images captured from Unmanned Aerial Vehicle (UAV) perspectives.
-We carefully curate SkyScenes images from Carla to comprehensively capture diversity across layout (urban and rural maps), weather conditions, times of day, pitch angles and altitudes with corresponding semantic, instance and depth annotations.
-Through our experiments using SkyScenes, we show that (1) Models trained on SkyScenes generalize well to different real-world scenarios, (2) augmenting training on real images with SkyScenes data can improve real-world performance, (3) controlled variations in SkyScenes can offer insights into how models respond to changes in viewpoint conditions, and (4) incorporating additional sensor modalities (depth) can improve aerial scene understanding.
-| Dataset | Reproducibility | Diversity | Annotations | Image Capture | Scale | |||||||
-| Metadata | Contr. Var. | Town | Daytime | Weather | Semantic | Instance | Depth | Altitude | Perspective | Resolution | ||
-| Real | ||||||||||||
-| 1 UAVid [19] | ✗ | ✗ | ✓ | ✗ | ✗ | ✓ | ✗ | ✗ | Med | Obl. | K | |
-| 2 AeroScapes [22] | ✗ | ✗ | ✗ | ✗ | ✗ | ✓ | ✗ | ✗ | (Low, Med) | (Obl., Nad.) | K | |
-| 3 ICG Drone [14] | ✗ | ✗ | ✗ | ✗ | ✗ | ✓ | ✗ | ✗ | Low | Nad. | K | |
-| Synthetic | ||||||||||||
-| 4 MidAir [9] | ✓ | partial | ✓ | ✓ | ✓ | ✓ | ✗ | ✓ | Low | (Obl., Nad.) | K | |
-| 5 VALID [2] | ✗ | ✗ | ✓ | ✓ | ✓ | ✓ | ✓ | ✓ | (Low, Med, High) | Nad. | K | |
-| 6 Espada [17] | ✗ | ✗ | ✓ | ✗ | ✗ | ✗ | ✗ | ✓ | (Med, High) | Nad. | K | |
-| 7 TartanAir [35] | ✓ | * | ✓ | ✓ | ✓ | ✓ | ✗ | ✓ | Low | (Fwd., Obl.) | M | |
-| 8 UrbanScene3D [16] | * | * | ✓ | * | * | ✗ | ✗ | ✗ | Med | Obl. | K | |
-| 9 SynthAer [30] | ✓ | ✓ | ✗ | ✓ | ✗ | ✓ | ✗ | ✗ | (Low, Med) | Obl. | K | |
-| 10 SynDrone [25] | ✓ | ✗ | ✓ | ✗ | ✗ | ✓ | ✗ | ✓ | (Low, Med, High) | (Obl., Nad.) | K | |
-| 11 SkyScenes | ✓ | ✓ | ✓ | ✓ | ✓ | ✓ | ✓ | ✓ | (Low, Med, High) | (Fwd., Obl., Nad.) | K | |
+We carefully curate SkyScenes images from Carla to comprehensively capture diversity across layouts (urban and rural maps), weather conditions, times of day, pitch angles and altitudes with corresponding semantic, instance and depth annotations.
+Through our experiments using SkyScenes, we show that (1) models trained on SkyScenes generalize well to different real-world scenarios, (2) augmenting training on real images with SkyScenes data can improve real-world performance, (3) controlled variations in SkyScenes can offer insights into how models respond to changes in viewpoint conditions (height and pitch), weather and time of day, and (4) incorporating additional sensor modalities (depth) can improve aerial scene understanding.
+Our dataset and associated generation code are publicly available at: https://hoffman-group.github.io/SkyScenes/
+Keywords:
+Aerial scene understanding, Synthetic-to-Real generalization, Segmentation, Domain Generalization, Synthetic Data
 1 Introduction
-We introduce SkyScenes, a densely annotated dataset of synthetic scenes captured from aerial (UAV) viewpoints under diverse layout (urban and rural), weather, daytime, pitch and altitude conditions.
-The setting of outdoor aerial imagery provides unique challenges for scene understanding: variability in altitude and angle of image capture, skewed representation for classes with smaller object sizes (humans, vehicles), size and occlusion variations of object classes in the same image, changes in weather or daytime conditions, etc.
+Aerial imagery provides a unique perspective that is invaluable for a wide range of applications, including surveillance [33, 39], mapping [37, 3], urban planning [24, 13], environmental monitoring [7, 29], and disaster response [31, 21].
+These applications rely on accurate and detailed analysis of aerial images to make informed decisions and effectively address various challenges.
 Naturally, training effective aerial scene-understanding models requires access to large-scale annotated exemplar data that have been carefully curated under diverse conditions.
@@
 However, carefully curating and annotating such images in the real-world can be prohibitively expensive due to various reasons.
-First, densely annotating high-resolution aerial images is expensive – for instance, densely annotating a single K image in UAVid [19] can take up to hours!
-Second, any effort to expand the real set to include widespread variations (weather, time of day, pitch, altitude) would be uncontrolled (i.e., can’t guarantee the same viewpoint under different conditions as real world is not static) and additionally would require re-annotating newly captured frames.
-Synthetic data curated from simulators can help counter both of these issues as – (1) labels are automatic and cheap to obtain and (2) it is possible to recreate the same viewpoint (with the same actor instances – vehicles, humans, etc.
+First, densely annotating every pixel of high-resolution real-world aerial images is expensive – for instance, densely annotating a single K image in UAVid [27] can take up to hours!
+Second, although diversity in training data is vital for developing robust generalization algorithms and for model sensitivity assessment, expanding the real set to include widespread variations (weather, time of day, pitch, altitude) would be uncontrolled (i.e., we can’t guarantee the same viewpoint under different conditions as the real world is not static), and hence would require re-annotating newly captured frames.
+Synthetic data curated from simulators can help counter both these issues as (1) labels are automatic and cheap to obtain and (2) it is possible to recreate same viewpoint (with same scene layout and actor instances, vehicles, humans, etc.
 in the scene) under differing conditions.
+The unique challenges introduced by outdoor aerial imagery setting such as variability in altitude and angle of image capture, skewed representation for classes with smaller object sizes (humans, vehicles), size and occlusion variations of object classes in the same image, etc., make it a relatively difficult problem compared to ground-view imagery setting.
+This is illustrated by the observation that the current developments in syn-to-real generalization methods [54, 16] have resulted in a significant reduction in the syn-to-real generalization gap in ground-view settings to as minimal as 20%111metric of choice: mIoU., compared to the comparatively higher margin observed in aerial imagery scenarios, reaching up to 50% (see Sec.
+0.D.3, 0.D.4, 0.D.5 in appendix).
 Unlike synthetic ground plane view datasets (especially for autonomous driving [46, 8, 56, 32, 43]), synthetic datasets for aerial imagery (see Table.
-1, rows -) relatively have received less attention [9, 2, 17, 35, 16, 30, 25] Existing synthetic datasets for aerial imagery can be lacking in a few different aspects – complementary metadata to reproduce existing frame viewpoints under different conditions, limited diversity, availability of dense annotations for a wide vocabulary of classes and image capture (height, pitch) conditions (see Table.
+1, rows -) have received relatively less attention [12, 5, 25, 53, 23, 47, 42].
+While existing synthetic aerial imagery datasets have tried to close this gap, they are often found lacking in a few different aspects – complementary metadata to reproduce existing frame viewpoints under different conditions, limited diversity, availability of dense annotations for a wide vocabulary of classes and image capture (height, pitch) conditions (see Table.
 1 for an exhaustive summary).
-We cover all these aspects by introducing SkyScenes, a synthetic dataset containing K densely annotated aerial scenes captured from Carla [7].
-We curate SkyScenes images by re-purposing the Carla [7] simulator for aerial viewpoints by re-positioning the agent camera to a desired altitude and pitch to obtain an oblique perspective.
-SkyScenes images are coupled with dense semantic, instance segmentation ( classes) and depth annotations.
-We carefully curate SkyScenes images by procedurally teleoperating the aerially situated camera through distinct map layouts across different weather and daytime conditions each over a combination of altitude and pitch variations (see Fig.
+These synthetic aerial datasets rarely allow for reproducing the exact viewpoint with different variations based on detailed scene metadata (see Table.1 Controlled Variations column), a key aspect for evaluating deep learning models’ responses to changing conditions and assessing sensitivity to single-variable alterations (e.g., weather, time of day, sensor angle).
+We cover all these aspects by introducing SkyScenes, a synthetic dataset containing k densely annotated aerial scenes, leveraging the Carla [10] simulator to capture diverse layout (urban and rural), weather, daytime, pitch and altitude conditions.
+Our pipeline meticulously generates scenes with precise actor locations and orientations, ensuring each scene’s reproducibility with a comprehensive metadata store and rigorous consistency checks.
+SkyScenes encompasses detailed semantic, instance segmentation ( classes), and depth annotations across distinct map layouts across different weather and daytime conditions each over a combination of altitude and pitch variations (see Fig.
 1 for examples).
@@
 Second, we store all metadata associated with the position of actors, camera, and other scene elements to be able to reproduce the same viewpoints under different weather and daytime conditions.
-Thirdly, we ensure that the generated data is physically realistic.
-This involves introducing variations in sensor locations, such as adding jitter to specified height and pitch values, to mimic real-world conditions.111Moreover, through rigorous validations, we ensure this process is consistent and yields error-free re-generations.
+Thirdly, we ensure that the generated data mimics real-world imperfections by introducing variations in sensor locations, such as adding jitter to specified height and pitch values.222Moreover, through rigorous validations, we ensure this process is consistent and yields error-free re-generations.
 See Sec.
-3.1 Finally, since Carla [7] by default does not spawn a lot of pedestrians in a scene, we propose an algorithm to ensure adequate representation of humans in the scene while curating images (see Sec.
+3.1.
+Finally, since Carla [10] by default does not spawn a lot of pedestrians in a scene, we propose an algorithm to ensure adequate representation of humans in the scene while curating images (see Sec.
 3.1 and Sec.
 3.2 for a detailed discussion).
+Our experiments across different real datasets and increasingly competitive semantic segmentation architectures consistently demonstrate that SkyScenes outperforms its closest synthetic counterpart dataset, SynDrone [42].
+However, despite these dataset-level improvements, the syn-to-real generalization gap persists, indicating that algorithmic improvements developed for ground-view imagery fail to translate effectively to aerial imagery.
+This underscores the urgent need for specialized algorithmic development in this area.
 Empirically, we demonstrate the utility of SkyScenes in several different ways.
-First, we show that SkyScenes is a “good” pre-training dataset for real-world aerial scene understanding by – (1) demonstrating that models trained on SkyScenes generalize well to multiple real-world datasets and (2) that SkyScenes pretraining can help reduce data requirements from the real-world (improved real-world performance in low-shot regimes).
+First, we show that SkyScenes is a valuable pre-training dataset for real-world aerial scene understanding by, (1) demonstrating that models trained on SkyScenes generalize well to multiple real-world datasets and (2) demonstrating that SkyScenes pretraining improves real-world performance in low-shot regimes.
 Second, we show that controlled variations in SkyScenes can serve as a diagnostic test-bed to assess model sensitivity to weather, daytime, pitch, altitude, and layout conditions – by testing SkyScenes trained models in unseen SkyScenes conditions.
@@
 To summarize, we make the following contributions:
-We introduce, SkyScenes, a densely-annotated dataset of synthetic aerial images.
+We introduce, SkyScenes, a densely-annotated dataset of k synthetic aerial images.
 SkyScenes contains images from different altitude and pitch settings, encompassing different layouts, weather, and daytime conditions with corresponding dense annotations and viewpoint metadata.
 We demonstrate that SkyScenes pre-trained models generalize well to real-world scenes and that SkyScenes data can effectively augment real-world training data for improved performance.
-We show that SkyScenes alone can serve as a diagnostic test-bed to assess model sensitivity to changing weather, daytime, pitch, altitude and layout conditions.
-Finally, we show that incorporating additional modalities (depth) while training aerial scene-understanding models can improve aerial scene recognition, enabling development of multi-modal segmentation models.
+We also bring attention to the point that while the synthetic-to-real gap has considerably narrowed for ground-view datasets, the same algorithms are unable to bridge this gap in aerial imagery.
+We show that our unique ability to generate controlled variations enables SkyScenes to serve as a diagnostic test-bed to assess model sensitivity to changing weather, daytime, pitch, altitude, and layout conditions.
+Finally, we show that incorporating additional modalities (depth) while training aerial scene-understanding models can improve aerial scene recognition, enabling further development of multi-modal segmentation models.
 2 Related Work
+| Dataset | Controlled Variations | Diversity | Annotation Diversity | Altitude | Perspective | Resolution | Scale | |||||
+| Town | Daytime | Weather | ||||||||||
+| Real | ||||||||||||
+| 1 UAVid [27] | ✗ | ✗ | ✗ | ✗ | S | Med | Obl. | k | ||||
+| 2 AeroScapes [34] | ✗ | ✗ | ✗ | ✗ | S | (Low, Med) | (Obl., Nad.) | k | ||||
+| 3 ICG Drone [20] | ✗ | ✗ | ✗ | ✗ | S | Low | Nad. | k | ||||
+| Synthetic | ||||||||||||
+| 4 Espada [25] | ✗ | ✓ | ✗ | ✗ | D | (Med, High) | Nad. | k | ||||
+| 5 UrbanScene3D [23] | ✗ | ✓ | ✗ | ✗ | - | Med | Obl. | k | ||||
+| 6 SynthAer [47] | ✓ | ✗ | ✓ | ✗ | S | (Low, Med) | Obl. | k | ||||
+| 7 MidAir [12] | ✗ | ✓ | ✓ | ✓ | (S,D) | Low | (Obl., Nad.) | k | ||||
+| 8 TartanAir [53] | ✗ | ✓ | ✓ | ✓ | (S,D) | Low | (Fwd., Obl.) | M | ||||
+| 9 VALID [5] | ✗ | ✓ | ✓ | ✓ | (S,I,D) | (Low, Med, High) | Nad. | k | ||||
+| 10 SynDrone [42] | ✗ | ✓ | ✗ | ✗ | (S,D) | (Low, Med, High) | (Obl., Nad.) | k | ||||
+| 11 SkyScenes | ✓ | ✓ | ✓ | ✓ | (S,I,D) | (Low, Med, High) | (Fwd., Obl., Nad.) | k | ||||
 Ground-view Synthetic Datasets.
-Real world ground-view scene-understanding datasets (Cityscapes [5], Mapillary [21], BDD-100K [37], Dark Zurich [28]) fail to capture the full range of variations that exist in the world.
+Real-world ground-view scene-understanding datasets (Cityscapes [8], Mapillary [32], BDD-100K [56], Dark Zurich [45]) fail to capture the full range of variations that exist in the world.
 Synthetic data is a popular alternative for generating diverse and bountiful views.
-GTAV [29], Synthia [26] and VisDA-C [23] are some of the widely-used synthetic datasets.
-These datasets can be curated using underlying simulators, such as the GTAV [29] game engine and Carla [7] simulator, and offer a cost-effective and scalable way to generate large amounts of labeled data under diverse conditions.
+GTAV [46], Synthia [43], and VisDA-C [38] are some of the widely-used synthetic datasets.
+These datasets can be curated using underlying simulators, such as GTAV [46] game engine and Carla [10] simulator and offer a cost-effective and scalable way to generate large amounts of labeled data under diverse conditions.
 Similar to SELMA [50] and SHIFT [49], we use Carla [10] as the underlying simulator for SkyScenes.
@@
 However, this study specifically focuses on lower altitudes, which are more relevant to UAVs, enabling object identification.
-Unfortunately, there is a scarcity of high-resolution real-world datasets based on UAV imagery that emphasize on object identification.
-Many existing urban scene datasets, like Aeroscapes [22], UAVid [19], VDD [1], UDD [4], UAVDT [8], VisDrone [38], Semantic Drones [14] and others, suffer from limited sizes and a lack of diverse images under different conditions.
+Unfortunately, there is a scarcity of high-resolution real-world datasets based on UAV imagery emphasizing object identification.
+Existing urban scene datasets, like Aeroscapes [34], UAVid [27], VDD [1], UDD [6], UAVDT [11], VisDrone [57], Semantic Drones [20] and others, suffer from limited sizes and a lack of diverse images under different conditions.
 This limitation raises concerns regarding model robustness and generalization.
@@
 Simulators can facilitate affordable, reliable, and quick collection of large synthetic aerial datasets, which aids in fast prototyping, improves real-world performance by enhancing robustness, and enables controlled studies on varied conditions.
-One such high-fidelity simulator, AirSim [31], used for the development and testing of autonomous systems (in particular, aerial vehicles), is the foundation of several synthetic UAV-based datasets like MidAir [9], Espada [17], Tartan Air [35], UrbanScene3d [16] and VALID [2].
+One such high-fidelity simulator, AirSim [48], used for development and testing of autonomous systems (in particular, aerial vehicles), is the foundation of several synthetic UAV-based datasets – MidAir [12], Espada [25], Tartan Air [53], UrbanScene3D [23] and VALID [5].
 Carla [10] is another such open-source simulator that is the foundation of datasets like SynDrone[42].
@@
 3 SkyScenes
-We curate SkyScenes using, Carla [7]222https://carla.org/ , which is a flexible and realistic open-source autonomous vehicle simulator.
+We curate SkyScenes using, Carla [10]333https://carla.org/ , which is a flexible and realistic open-source autonomous vehicle simulator.
 The simulator offers a wide range of sensors, environmental configurations, and varying rendering configurations.
 As noted earlier, we take several important considerations into account while curating SkyScenes images.
-These include strategies for obtaining diverse synthetic data and embedding real-world irregularities, avoiding correlated images, addressing skewed class representations and more.
-In this section, we first discuss such desiderata, and then describe our procedural image curation algorithm.
+These include strategies for obtaining diverse synthetic data and embedding real-world irregularities, avoiding correlated images, addressing skewed class representations, and more.
+In this section, we first discuss such desiderata and then describe our procedural image curation algorithm.
 Finally, we describe different aspects of the curated dataset.
 3.1 (Synthetic) Aerial Image Desiderata
-Before delving into the image curation pipeline, we first outline a set of desiderata taken into account while curating synthetic aerial images in SkyScenes.
+Before investigating the image curation pipeline, we first outline a set of desiderata taken into account while curating synthetic aerial images in SkyScenes.
 1.
+Viewpoint Reproducibility: Critical to understanding how models respond to changing conditions is the ability to evaluate them under scenarios where only one variable is altered.
+However, any effort to do so in the real-world would be uncontrolled, due to its dynamic (constantly changing) nature.
+In contrast, simulated data allows us to do so by providing control over image generation conditions.
+Unlike certain existing aerial datasets that do not support this feature (see Table.
+1), we do so in SkyScenes by additionally storing comprehensive metadata for each viewpoint (and image), including details about camera world coordinates, orientation, and all movable/immovable actors and objects in the scene.
+We couple this with rigorous consistency checks for image generation that verify the number of actors, their location, sensor height, pitch, etc.
+This meticulous approach enables us to reproduce the same viewpoint under multiple conditions effortlessly.
+2 .
+Adequate Representation of Tail Classes: Unlike ground-view datasets, pixel distribution of classes in aerial images is substantially more long-tailed (see Fig.
+2 (a); classes with smaller object size, humans).
+This substantial difference in class proportions severely affects the performance of tail classes in aerial datasets when compared to ground-view datasets (see Fig.
+2 (b)), thus making visual recognition tasks harder.
+To counter this, we consider structured spawning of humans to ensure adequate representation (see Sec.
+3.2).
+3.
 Adequate Height Variations: Aerial images are captured at different altitudes to meet specific needs.
@@
 With SkyScenes, our aim is to provide flexibility in altitude sampling, thus accommodating various real-world requirements.
-We curate SkyScenes images at heights of m, m and m.
-Additionally, recognizing imperfections in real-world actuation, we induce slight jitter in the height values (jitter ) to simulate realistic data sampling.
-2.
+We curate SkyScenes images at heights of m, m, and m.
+Additionally, recognizing imperfections in real-world actuation, we induce slight jitter in the height values (m) to simulate realistic data sampling.
+4.
 Adequate Pitch Variations: Similar to height, aerial images can be captured from primary perspectives or pitch angles (): nadir (), oblique (), or forward () views (see Table.
@@
 Oblique views, on the other hand, capture objects from a side profile, aiding object recognition and providing valuable context and depth perspective often lost in nadir and forward views.
-To ensure widespread utility, SkyScenes data generation process is designed to support all these viewing angles, with a particular emphasis on oblique views (most common one).
-Similar to height, pitch variations allow models trained on SkyScenes to generalize to different real-world conditions.
-We use for oblique-views and introduce jitter (jitter ) to mimic real-world data sampling.
-3.
+To ensure widespread utility, SkyScenes data generation process is designed to support all these viewing angles, with a particular emphasis on oblique views (the most common one).
+Similar to height, pitch variations allow models trained on SkyScenes to generalize to different viewpoint variations.
+We use for oblique-views and introduce jitter () to mimic real-world data sampling.
+5.
 Adequate Map Variations: In addition to sensor locations, it is equally important to curate aerial images across diverse scene layouts.
@@
 4 illustrates how images curated from different towns in Carla [10] differ in class distributions.
-4.
+6.
 Adequate Weather & Daytime Variations: Training robust perception models using SkyScenes that generalize to unforeseen environmental conditions necessitates the curation of annotated images encompassing various weather and daytime scenarios.
-To accomplish this, we generate SkyScenes images from identical viewpoints under different variations – ClearNoon, ClearSunset, MidRainNoon, ClearNight and CloudyNoon.333Note that Carla [7] provides such conditions but we use only such conditions in this preliminary version of SkyScenes.
+To accomplish this, we generate SkyScenes images from identical viewpoints under different variations – ClearNoon, ClearSunset, MidRainNoon, ClearNight, and CloudyNoon.444Note that Carla [10] provides such conditions but we use only such conditions in this preliminary version of SkyScenes.
 Generating images in different conditions from the same perspectives allows us to (1) leverage diverse data for improved generalization and (2) systematically investigate the susceptibility of trained models to variations in daytime and weather conditions.
-5.
+7.
 Fine-grained Annotations: To support a host of different computer vision tasks (segmentation, detection, multimodal recognition), we curate all SkyScenes images with dense semantic, instance segmentation and depth annotations.
@@
 1 column for an example).
-Figure 2: Ground-View vs Aerial-View Pixel Proportions.
-For a subset of commonly annotated classes across Cityscapes [5] (red), UAVid [19] (dark blue) and SkyScenes (light blue), we show the percentage of pixels occupied by different classes.
-Aerial scenes (in UAVid) have significant under-representation of tail classes (vehicle, human).
-SkyScenes image curation (for the same viewpoint settings as UAVid) helps counter this discrepancy via synthetic aerial scenes.
--
-6.
-Viewpoint Reproducibility: Critical to understanding how models respond to changing conditions is the ability to evaluate them under scenarios where only one variable is altered.
-However, any effort to do so in the real-world would be uncontrolled, due to it’s dynamic (constantly changing) nature.
-In contrast, simulated data allows us to do so by providing control over image generation conditions.
-Unlike certain existing aerial datasets that do not support this feature (see Table.
-1), we do so in SkyScenes by additionally storing comprehensive metadata for each viewpoint (and image), including details about camera world coordinates, orientation and all movable / immovable actors and objects in the scene.
-We couple this with rigorous consistency checks for image generation that verify the number of actors, their location, sensor height, pitch, etc.
-This meticulous approach enables us to reproduce the same viewpoint under multiple conditions effortlessly.
-7 .
-Adequate Representation of Tail Classes: Unlike ground-view datasets, pixel distribution of classes in aerial images is substantially more long-tailed (see Fig.
-2; classes with smaller object size, humans), making visual recognition tasks harder.
-This problem is further exacerbated by differing class distributions across synthetic and real data (see Fig.
-5) – especially severe for instances of human class.
-To counter this, we consider structured spawning of humans to ensure adequate representation.
 3.2 SkyScenes Image Generation
@@
 We accomplish the first by mimicking a UAV perspective in Carla [10] by positioning the ego vehicle (with RGB, semantic and depth sensors) based on specified (high) altitude () and pitch () values to generate aerial views (see Fig.
-3).444This also requires setting other scene – weather, daytime, etc.
-– and camera – notably the (field of view) and (image resolution) – parameters.
-Once positioned, the agent is translated by fixed amounts to traverse the scene and capture images from various viewpoints (detailed in Algo.
-1 in the appendix).
-Initially, we generate datapoints for each of the town variations under ClearNoon conditions using the baseline setting.
+2(a)).555This also requires setting other scenes – weather, daytime, etc.
+– and camera (notably the (field of view) and image resolution x) parameters.
+Once positioned, the agent is translated by fixed amounts to traverse the scene and capture images from various viewpoints (detailed in Sec.
+0.B.1 in the appendix).
+Initially, we generate data points for each of the town variations under ClearNoon conditions using the baseline m setting.
 Subsequently, following
-the traversal algorithm (see Algo. 1 in the appendix),
+the traversal algorithm (see Sec. 0.B.1 in the appendix),
 we re-generate these datapoints across weather conditions and height/pitch variations,
 resulting in images.
+| mIoU | ||
+| Eval Data | human | All |
+| SkyScenes | ||
+| 1 Before HS | ||
+| 2 After HS | 61.79 | 84.07 |
+| SkyScenes UAVid | ||
+| 3 Before HS | ||
+| 4 After HS | 10.21 | 47.09 |
 Checks and Balances.
@@
 This also helps with moving dynamic actors by a considerable amount in the scene.
-Additionally, pedestrian objects are regenerated before saving an image, which adds randomness to the spawning and placement of pedestrians and further reduces correlation between frames.
-| Layout | RS (%) | HS (%) |
-| 1 Town01 | ||
-| 2 Town02 | ||
-| 3 Town03 | ||
-| 4 Town04 | ||
-| 5 Town05 | ||
-| 6 Town06 | ||
-| 7 Town07 | ||
-| 8 Town10HD | ||
-| 9 All | ||
-| 10 UAVid [19] |
-| mIoU | |||
-| Eval Data | HS | human | All |
-| 1 S | |||
-| 2 S | ✓ | 61.79 | 84.07 |
-| 3 SU | |||
-| 4 SU | ✓ | 10.21 | 47.09 |
+Additionally, pedestrian objects are regenerated before saving an image, which adds randomness to the spawning and placement of pedestrians, reducing the correlation between frames.
 Adequate Representation of humans.
 Real-world scenes often exhibit a long-tailed distribution in pixel proportions, particularly in aerial images where variations in object sizes and camera positions contribute to significant under-representation of the tail classes (in Fig.
-2, for the shared set of classes across UAVid [19] (aerial) and Cityscapes [5] (ground), we can see that the class distributions are different and aerial images are significantly more heavy tailed).
+2, for the shared set of classes across UAVid [27] (aerial) and Cityscapes [8] (ground), we can see that the class distributions are different and aerial images are significantly more heavy-tailed).
 As a result, naively spawning humans (rarest class) in Carla [10] is detrimental for eventual task performance – for the human class, a SkyScenes trained DAFormer [16] (with HRDA [17] source training; MiT-B5 [55] backbone) model leads to an in-distribution performance of mIoU and out-of-distribution (SkyScenes UAVid [27]) performance of mIoU.
-To counter this under-representation issue, we design an algorithm, HumanSpawn()555More details in Sec.
-B.1 of the appendix., to explicitly spawn more human instances while curating SkyScenes images.
-HumanSpawn() increases human instances by per snapshot, improving the proportion of densely annotated humans in SkyScenes by approximately times (see Table.
-3 & Fig.
-4).
+To counter this under-representation issue, we design an algorithm, HumanSpawn() (see Sec.
+0.B.1 in appendix), to explicitly spawn more human instances while curating SkyScenes images.
+HumanSpawn() increases human instances by per snapshot, improving the proportion of densely annotated humans in SkyScenes by approximately times (see Fig.
+3 (a) & Fig.
+3 (b)).
 This improvement in human representation is also evident in eventual task performance, with in-distribution and out-of-distribution mIoUs for humans increasing from to () and to () respectively (see Table.
-3.)
+3 (c)).
 3.3 SkyScenes: Dataset Details
@@
 These are building, fence, pedestrian, pole, roadline (markings on road), road, sidewalk, vegetation, cars, wall, traffic sign, sky, bridge, railtrack, guardrail, traffic light, water, terrain, rider, bicycle, motorcycle, bus, truck and others (see Fig.
-1 for an example).666We provide detailed definitions in Sec.
-B.2 of the appendix.
+1 for an example and Sec.
+0.B.2 in appendix for definitions)
 Training, Validation and Test Splits.
 SkyScenes has images per town (across towns) for each of the weather and daytime conditions, and height & pitch combinations, resulting in a total of images.
-We use % ( images) of the dataset for training models, with % ( images) each for validation and testing.
-While creating train, val and test splits, we collect equal number of samples from each town by dividing each town-specific traversal sequence into segments: the initial % for training, the next % for testing, and the final % of the segment for validation.
-Moreover, within each split, we ensure that every viewpoint is accompanied by its different variations across weather, daytime, height, and pitch settings.
-This safeguards against any potential cross-contamination across different splits while ensuring fair representation and equal distributions of all variations.
+We use % ( images) of the dataset for training models, with % ( images) each for validation and testing (see Sec.
+0.C.2 in appendix).
 Class Distribution(s).
 In Fig.
-6, we compare the number of densely annotated pixels in SkyScenes with those in Syndrone [25] (another synthetic dataset) and UAVid [19] (a real aerial dataset).
-Compared to both the real and prior synthetic counterparts, we show that SkyScenes is specifically curated to ensure adequate tail class representation (more human and vehicle pixels) to facilitate model learning.
-Additionally, in Fig.
-5, we highlight how the distribution of classes changes across variations within SkyScenes– rural and urban map layouts and height and pitch specifications.
-SkyScenes exhibits substantial diversity in class-distributions across such conditions, allowing these individual conditions to serve as diagnostic splits to assess model sensitivity (see Sec.
-4.2)
+4, we highlight how the distribution of classes changes across variations within SkyScenes– rural and urban map layouts and height and pitch specifications.
+SkyScenes exhibits substantial diversity in class distributions across such conditions, allowing these individual conditions to serve as diagnostic splits to assess model sensitivity (see Sec.
+4.2).
 4 Experiments
-We conduct semantic segmentation experiments with SkyScenes to assess a few different factors.
-First, we check if training on SkyScenes is beneficial for real-world transfer.
-Second, we check if SkyScenes can augment real-world training data in low and full shot regimes.
-Third, we check if variations in SkyScenes can be used to assess sensitivity of trained models to changing conditions.
-Finally, we check if using additional modality information (depth) can help improve aerial scene understanding.
-Synthetic and Real Datasets.
-We compare real-world generalization performance of training on SkyScenes with SynDrone [25], a recently proposed synthetic aerial dataset also curated from Carla [7].
-We assess performance on real-world aerial datasets – UAVid [19], AeroScapes [22], ICG Drone [14].
-Since different datasets have different class vocabularies and definitions, for our experiments, we adapt the class vocabulary of the synthetic source dataset to that of the target real-world datasets.777We detail class merging and assignment schemes used for these experiments in Table.
-8, Table.
-9 and Table.
-10 in the appendix.
-Additionally, since different real aerial datasets have been captured from different heights and pitch angles, we train models on subsets of synthetic datasets that are aligned with corresponding real data conditions.
-We provide additional details for the real aligned synthetic data selection and model evaluation in Sec.
-D.1 of the appendix.
-Models.
-We use both CNN – DeepLabv2 [3] (ResNet-101 [10]) – and transformer – DAFormer [11] (with HRDA [12] source training; MiT-B5 [36] backbone) – based semantic segmentation architectures for our experiments.
-We provide implementation details surrounding our experiments in Sec.
-C of the appendix.
 | Source | (Target) Real-World mIoU () | ||
@@
 | 2 SkyScenes | 41.82 | 26.94 | 15.14 |
-| 3 Train-on-Target | |||
 | DAFormer (MiT-B5) [17] | |||
-| 4 SynDrone | |||
-| 5 SkyScenes | 47.09 | 40.72 | 25.91 |
-| 6 Train-on-Target | |||
-| Source | (Target) Real-World mIoU () | |||||
-| UAVid | AeroScapes | ICG Drone | ||||
-| vehicle | human | vehicle | person | vehicle | person | |
-| 1 SynDrone | ||||||
-| 2 SkyScenes | 63.64 | 10.21 | 80.99 | 3.09 | 39.71 | 45.89 |
-| 3 Train-on-Target | ||||||
+| 3 SynDrone | |||
+| 4 SkyScenes | 47.09 | 40.72 | 25.91 |
+| Rein (DINOv2) | |||
+| 5 SynDrone | 54.92 | ||
+| 6 SkyScenes | 43.96 | 28.10 | |
+| Source | (Target) Real-World IoU () | |||||||
+| UAVid | AeroScapes | ICG Drone | ||||||
+| vehicle | human | vehicle | person | vehicle | person | |||
+| DAFormer (MiT-B5) [17] | ||||||||
+| 1 SynDrone | ||||||||
+| 2 SkyScenes | 63.64 | 10.21 | 80.99 | 3.09 | 39.71 | 45.89 | ||
+| Rein (DINOv2) [54] | ||||||||
+| 3 SynDrone | ||||||||
+| 4 SkyScenes | 75.14 | 25.52 | 87.71 | 21.67 | 50.91 | 77.93 | ||
+We conduct semantic segmentation experiments with SkyScenes to assess a few different factors.
+First, we check if training on SkyScenes is beneficial for real-world transfer.
+Second, we check if SkyScenes can augment real-world training data in low and full shot regimes.
+Third, we check if variations in SkyScenes can be used to assess the sensitivity of trained models to changing conditions.
+Finally, we check if using additional modality information (depth) can help improve aerial scene understanding.
+Synthetic and Real Datasets.
+We compare real-world generalization performance of training on SkyScenes with SynDrone [42], a recently proposed synthetic aerial dataset also curated from Carla [10] featuring different conditions across different map layouts.
+We assess performance on real-world aerial datasets – UAVid [27], AeroScapes [34], ICG Drone [20].
+Since different datasets have different class vocabularies and definitions, for our experiments, we adapt the class vocabulary of the synthetic source dataset to that of the target real-world datasets (see Sec.
+0.C.1 in appendix for class merging and assignment schemes).
+Additionally, since different real aerial datasets have been captured from different heights and pitch angles, we train models on subsets of synthetic datasets that are aligned with corresponding real data conditions.
+We provide additional details for the real aligned synthetic data selection and model evaluation in Sec.
+0.D.3 in the appendix.
+Models.
+We use (1) CNN – DeepLabv2 [4] (ResNet-101 [14]), (2) transformer – DAFormer [16] (with HRDA [17] source training; MiT-B5 [55] backbone) and (3) Vision Foundation Model – Rein [54] (LoRA [19] fine-tuned Dino-V2 [36] backbone) based semantic segmentation architectures for our experiments.
+We provide implementation details for our experiments in Sec.
+0.C in appendix.
+Figure 5: SkyScenes RGB Images and corresponding depth images generated using depth sensor for , , and ClearNoon setting across four different town layouts.
+Sensors
+SkyScenes Test IoU ()
+clutter
+building
+road
+tree
+low-veg.
+vehicle
+human
+Avg
+1 RGB
+2 RGB+D
+90.64
+95.97
+94.87
+89.41
+74.36
+86.87
+50.47
+83.22
+Table 5: Multi-modal Segmentation in SkyScenes.
+We evaluate M3L [28] multimodal segmentation architectures with MiT-B5 [55] backbones using RGB and RGB+D data in SkyScenes.
+Additional sensors improve aerial scene understanding significantly across various classes in UAVid [27].
 4.1 SkyScenes Real Transfer
@@
 In Table.
-4, we show how models trained on SkyScenes exhibit strong out-of-the box generalization performance on multiple real world datasets.
-We find that SkyScenes pretraining exhibits stronger generalization compared to SynDrone [25] across both CNN and transformer segmentation backbones.
+3, we show how models trained on SkyScenes exhibit strong out-of-the box generalization performance on multiple real world datasets.
+We find that SkyScenes pretraining exhibits stronger generalization compared to SynDrone
+[42] across both CNN and transformer segmentation backbones.
 In Table.
-5, we show how generalization improvements are more pronounced for under-represented tail classes (vehicles and humans)888Model performance comparisons for all the classes are provided in Table 11, Table.
-12 and Table.
-13 in the appendix.
+3, we show how generalization improvements are more pronounced for under-represented tail classes (vehicles and humans).
+Comparison across all classes is provided in Tables 11, 12 and 13 in appendix.
 SkyScenes can augment real training data.
 In addition to zero-shot real-world generalization, akin to other synthetic aerial datasets, we also show how SkyScenes is useful as additional training data when labeled real-world data is available.
-In Fig.
-7, for SkyScenes UAVid [19], we compare models trained only using of the UAVid [19] training images with counterparts that were either pretrained using SkyScenes data or additionally supplemented with SkyScenes data at training time.
-We find that in low-shot regimes (when little “real” world data is available), SkyScenes data (either explicitly via joint training or implicitly via finetuning) is beneficial in improving recognition performance.
-We find this to be especially beneficial for under-represented classes in aerial imagery (such as humans and vehicles).
-We discuss more such results across other real-world datasets in Sec.
-D.2 of the appendix.
-| Train | Test mIoU () | ||
-| Clear | Cloudy | Rainy | |
-| 1 Clear | |||
-| 2 Cloudy | |||
-| 3 Rainy | |||
-| Train | Test mIoU () | ||
-| Noon | Sunset | Night | |
-| 1 Noon | |||
-| 2 Sunset | |||
-| 3 Night | |||
-| Train | Test mIoU () | |
-| Rural | Urban | |
-| 1 Rural | ||
-| 2 Urban | ||
-| Test mIoU () | ||||
-| Height | Pitch | |||
-| 1 m | 48.50 | |||
-| 2 m | ||||
-| 3 m | ||||
+In Table.
+4, for SkyScenes UAVid [27], we compare models trained only using , , of the UAVid [27] training images with counterparts that were either pretrained using SkyScenes data or additionally supplemented with SkyScenes data at training time.
+We find that in low-shot regimes (when little “real” world data is available), SkyScenes data (either explicitly via joint training or implicitly via finetuning) is beneficial in improving recognition performance (see Sec.
+0.D.5 of appendix).
 4.2 SkyScenes as a Diagnostic Framework
@@
 4 for variations in class distributions).
-We curate images under such diverse conditions in a controlled manner – ensuring the same spatial coordinates for variations, same spatial coordinates and settings across different weather and daytime conditions, same number of images across layouts.
+We curate images under such diverse conditions in a controlled manner – ensuring the same spatial coordinates for variations, same spatial coordinates and settings across different weather and daytime conditions, the same number of images across layouts.
 This allows us to assess the sensitivity of trained models to one factor of variation (, , daytime, weather, map layout) by changing that specific aspect.
@@
 In Table.
-6 (d), we evaluate a model trained under moderate conditions under different variations.
+6 (d), we evaluate a model trained under moderate m conditions under different variations.
 We find that as altitudes increase, trained models are better at recognizing objects from oblique () viewpoints.
 We provide exhaustive quantitative comparisons in Sec.
-D.3 of the appendix.
-| Sensors | SkyScenes Test mIoU () | |||||||
-| clutter | building | road | tree | low-veg. | vehicle | human | Avg | |
-| 1 RGB | ||||||||
-| 2 RGB+D | 90.64 | 95.97 | 94.87 | 89.41 | 74.36 | 86.87 | 50.47 | 83.22 |
-4.3 SkyScenes Enables Multi-modal Recognition
-Sensors on UAVs in deployable settings are not limited to RGB cameras.
-It is common to have UAVs deployed in the real-world with additional modality sensors, such as depth.
-Additional sensor modalities can also potentially help improve aerial scene understanding.
+20 in the appendix.
+4.3 SkyScenes Enables Multi-modal Dense Prediction
+Sensors on UAVs in deployable settings often include modalities beyond RGB cameras, such as depth sensors.
+These additional modalities can significantly enhance aerial scene understanding.
 In Table.
-7, we check if assisting RGB with Depth observations for SkyScenes viewpoints can help improve aerial semantic segmentation using M3L [20], a model capable of multimodal segmentation.
+5, we investigate the impact of augmenting RGB data with depth observations from SkyScenes viewpoints on aerial semantic segmentation using M3L [28], a multimodal segmentation model.
 Similar to our DAFormer [16] experiments, we consider a SegFormer equivalent version of M3L [28] (with an MiT-B5 [55] backbone).
 We test RGB and RGB+D models trained under (moderate viewpoint) conditions on SkyScenes and find that incorporating additional Depth observations can substantially improve recognition performance.
-This demonstrates that annotated images in SkyScenes can be used to train multimodal scene-recognition models.
+This demonstrates that images in SkyScenes can be used to train multimodal scene-recognition models.
 5 Conclusion
-We introduce SkyScenes, a large-scale densely-annotated dataset of synthetic aerial scenes curated from unmanned aerial vehicle (UAV) perspectives.
-We collect SkyScenes images from Carla by first aerially situating an agent (to get an aerial perspective) and then procedurally tele-operating the agent through the scene to capture aerial frames with corresponding semantic, instance and depth annotations.
-Our careful curation process ensures that SkyScenes images are carefully curated across diverse weather, daytime, map, height and pitch conditions, with accompanying metadata that enables reproducing the same viewpoint (spatial coordinates and perspective) under differing conditions.
-Through our experiments, we demonstrate how (1) SkyScenes trained models can generalize to real-world settings, (2) SkyScenes images can augment labeled real-world data in low-shot regimes, (3) SkyScenes can serve as a diagnostic framework to assess model sensitivity to changing conditions and (4) additional sensors, such as Depth, in SkyScenes can facilitate development of multi-modal aerial scene understanding models.
-Lastly, we plan on updating SkyScenes with evolving considerations for real-world aerial scene-understanding – improved realism, additional anticipated edge cases – as more and more features are supported in the underlying simulator.
-We intend to release both the dataset and associated generation code for SkyScenes publicly, and hope that our experimental findings encourage further research using SkyScenes for aerial scenes.
+We introduce SkyScenes, a large-scale densely-annotated dataset of synthetic aerial scene images curated from unmanned aerial vehicle (UAV) perspectives.
+SkyScenes images are generated using Carla by situating an agent aerially and procedurally tele-operating it through the scene to capture frames with semantic, instance, and depth annotations.
+Our careful curation process ensures that SkyScenes images span across diverse weather, daytime, map, height, and pitch conditions, with accompanying metadata that enables reproducing the same viewpoint (spatial coordinates and perspective) under differing conditions.
+Through our experiments, we demonstrate that: (1) SkyScenes-trained models generalize well to real-world settings, (2) SkyScenes augments labeled real-world data in low-shot scenarios, (3) SkyScenes serves as a diagnostic tool for assessing model sensitivity to varied conditions, and (4) incorporating additional sensors like depth enhances multi-modal aerial scene understanding.
+We aim to enhance SkyScenes with improved realism, additional anticipated edge cases, and support for 3D perception tasks aligning with advancements in our simulator (additional details in Sec.
+0.F of appendix) We have publicly released the dataset and associated generation code and hope that our experimental findings encourage further research using SkyScenes for aerial scenes.
 Acknowledgements.
@@
 This work has been partially sponsored by NASA University Leadership Initiative (ULI) #80NSSC20M0161, ARL, and NSF #2144194.
-Appendix A Appendix
-This appendix is organized as follows.
+Appendix 0.A Overview
+The appendix is organized as follows.
 In Sec.
-B, we provide more details on different aspects of dataset – including the procedural image curation algorithm, algorithm to ensure appropriate representation of human instances and brief descriptions of all the classes in SkyScenes.
+0.B.1, we provide more details on different aspects of dataset – including the procedural image curation algorithm, algorithm to ensure appropriate representation of human instances, brief descriptions of all the classes in SkyScenes and comparisons between data distribution of SkyScenes and real datasets.
 Then, we describe experimental details in Sec.
-C – class-merging schemes used for our SkyScenes Real transfer experiments, train / val / test splits, experimental details for SkyScenes diagnostic setup to probe model vulnerabilities.
+0.C – class-merging schemes used for our SkyScenes Real transfer experiments, train / val / test splits and experimental details for SkyScenes diagnostic setup to probe model vulnerabilities.
 Sec.
-D provides more quantitative and qualitative experimental results.
-Appendix B SkyScenes Details
-B.1 Image Generation Algorithms
+0.D provides more quantitative and qualitative experimental results.
+Appendix 0.B SkyScenes Details
+0.B.1 Image Generation Algorithms
 We curate SkyScenes using Carla [10] simulator.
@@
 This also requires setting other scene information like – town, weather, and daytime.
-The camera (field of view) and (image resolution) are also set.
+The camera (field of view) and = (image resolution) are also set.
 To maintain the adequate representation of tail classes, especially for humans we incorporate structured spawning of humans using Algo.
@@
 Manual spawning not only increases the number of human instances and their proportion but also aligns their placement with real-world settings.
+As SkyScenes images are curated by teleoperating over entire maps (rural or urban) across multiple layouts that differ substantially in class distributions (Fig.
+4 left in the paper).
+Since these frames are stored with corresponding geographical (layout identifier) and positional (spatial location) metadata, filtering data splits to avoid overlap across physical layout regions is always possible.
 The steps involved in manual spawning instances of humans are summarized below:
@@
 1), we re-generate these datapoints across weather conditions and height/pitch variations, resulting in images.
-B.2 Class Descriptions
+0.B.2 Class Descriptions
 We provide semantic, instance, and depth annotations for every image in SkyScenes. SkyScenes provides dense semantic annotations for classes. These are:
@@
 vegetation: trees, hedges, all kinds of vertical vegetation (ground-level vegetation is not included here).
-cars: cars in scene
+cars: cars
 wall: individual standing walls, not part of buildings
@@
 sky: open sky, including clouds and sun
-ground: any horizontal ground-level structures that does not match any other category
+ground: any horizontal ground-level structures that do not match any other category
 bridge: the structure of the bridge
@@
 truck: trucks in scenes
-B.3 Depth & Instance Segmentation
+0.B.3 Depth & Instance Segmentation
 We also provide depth and instance segmentation annotations along with semantic segmentation annotations.
@@
 We also provide depth-aided semantic segmentation results (training details in Sec.
-C.3).
-Appendix C Experiment Details
-C.1 Class Merging Details
-| Syn UAVid | UAVid | SkyScenes | SynDrone |
-| 1 clutter | clutter | unlabelled | unlabelled |
-| other | other | ||
-| traffic sign | traffic sign | ||
-| rail track | rail track | ||
-| guard rail | guard rail | ||
-| traffic light | traffic light | ||
-| static | static | ||
-| dynamic | dynamic | ||
-| ground | ground | ||
-| sidewalk | sidewalk | ||
-| fence | fence | ||
-| sky | sky | ||
-| bridge | bridge | ||
-| water | water | ||
-| pole | pole | ||
-| 2 building | building | building | building |
-| wall | wall | ||
-| 3 road | road | road | road |
-| roadline | roadline | ||
-| 4 vegetation | vegetation | vegetation | vegetation |
-| 5 low vegetation | terrain | terrain | terrain |
-| 6 person | person | pedestrian | pedestrian |
-| rider | rider | ||
-| motorcycle | motorcycle | ||
-| bicycle | bicycle | ||
-| 7 vehicle | car | car | car |
-| truck | truck | ||
-| bus | bus | ||
+0.C.4).
+0.B.4 SkyScenes vs Real Characteristics
+| FID Score () | |||
+| UAVid | AeroScapes | ICG Drone | |
+| 1 SynDrone | |||
+| 2 SkyScenes | 8.78 | 7.66 | 5.77 |
+In Fig.
+7, we compare perceptual similarity (in terms of FID [15]) of synthetic images from SkyScenes and SynDrone [42] with real datasets (UAVid [27], AeroScapes [34], ICG Drone [20]).
+We find SkyScenes images are closer to real data distributions than SynDrone [42].
+For semantic comparisons, we show class-distribution comparisons (per-class pixel frequencies) between SkyScenes, SynDrone [42] and UAVid [27] in Fig 7.
+SkyScenes images display higher distribution of tail classes(vehicles, humans) compared to SynDrone [42], and shows better alignment with the real data distribution (UAVid [27]).
+Appendix 0.C Experiment Details
+0.C.1 Class Merging Details
+| Syn UAVid | UAVid | SkyScenes | SynDrone | VALID | SynthAer |
+| 1 clutter | clutter | unlabelled | unlabelled | background | sky |
+| sidewalk | sidewalk | pavement | footpath | ||
+| fence | fence | fence | |||
+| bridge | bridge | bridge | |||
+| water | water | water | |||
+| traffic light | traffic light | traffic light | |||
+| other | other | sign | |||
+| traffic sign | traffic sign | land | |||
+| rail track | rail track | tunnel | |||
+| guard rail | guard rail | pool | |||
+| static | static | stones | |||
+| dynamic | dynamic | pierruble | |||
+| ground | ground | chair | |||
+| sky | sky | ice | |||
+| pole | pole | ship | |||
+| plane | |||||
+| harbor | |||||
+| lamp | |||||
+| bus stop | |||||
+| powerline | |||||
+| garbage bin | |||||
+| other low obstacle | |||||
+| other high obstacle | |||||
+| 2 building | building | building | building | building | building |
+| wall | wall | wall | |||
+| 3 road | road | road | road | road | road |
+| roadline | roadline | ||||
+| 4 vegetation | vegetation | vegetation | vegetation | tree | tree |
+| 5 low vegetation | terrain | terrain | terrain | other plant | vegetation |
+| 6 person | person | pedestrian | pedestrian | person | |
+| rider | rider | ||||
+| motorcycle | motorcycle | ||||
+| bicycle | bicycle | ||||
+| 7 vehicle | car | car | car | small vehicle | car |
+| truck | truck | large vehicle | |||
+| bus | bus | ||||
 | train |
+| Common Scheme | UAVid | AeroScapes | ICG Drone | SkyScenes |
+| 1 clutter | clutter | background | unlabeled | unlabeled |
+| sky | dirt | sky | ||
+| grass | pole | |||
+| gravel | sidewalk | |||
+| water | traffic sign | |||
+| pool | other | |||
+| rock | ground | |||
+| fence-pole | guard rail | |||
+| dog | traffic light | |||
+| ar-marker | static | |||
+| obstacle | dynamic | |||
+| conflicting | water | |||
+| terrain | ||||
+| 2 road | road | road | paved area | road |
+| roadline | ||||
+| rail track | ||||
+| 3 nature | vegetation | vegetation | vegetation | vegetation |
+| tree | tree | |||
+| bald tree | ||||
+| 4 person | human | person | person | pedestrian |
+| bicycle | bicycle | bicycle | ||
+| rider | ||||
+| motorcycle | ||||
+| 5 vehicle | car | vehicle | car | car |
+| truck | ||||
+| bus | ||||
+| 5 construction | building | building | roof | building |
+| wall | wall | |||
+| fence | fence | |||
+| window | bridge | |||
+| door |
 | Syn AeroScapes | AeroScapes | SkyScenes | SynDrone |
@@
 | 8 sky | sky | sky | sky |
-| Syn ICG Drone | ICG Drone | SkyScenes | SynDrone |
-| 1 other | obstacle | unlabelled | unlabelled |
-| dog | other | other | |
-| conflicting | traffic sign | traffic sign | |
-| ar-marker | sky | sky | |
-| unlabelled | bridge | bridge | |
-| gurad rail | guard rail | ||
-| traffic light | traffic light | ||
-| static | static | ||
-| dynamic | dynamic | ||
-| 2 fence | fence | fence | fence |
-| 3 pole | fence pole | pole | pole |
-| 4 vegetation | tree | vegetation | vegetation |
-| bald-tree | |||
-| 5 building | wall | wall | wall |
-| roof | building | building | |
-| door | |||
-| window | |||
-| 6 water | water | water | water |
-| pool | |||
-| 7 bicycle | bicycle | bicycle | bicycle |
-| motorcycle | motorcycle | ||
-| 8 vehicle | car | car | car |
-| truck | truck | ||
-| bus | bus | ||
-| train | |||
-| 9 person | person | pedestrian | pedestrian |
-| rider | rider | ||
-| 10 paved area | paved area | ground | ground |
-| sidewalk | sidewalk | ||
-| road | road | ||
-| roadline | roadline | ||
-| rail track | rail track | ||
-| 11 terrain | rocks | terrain | terrain |
-| gravel | |||
-| dirt | |||
-| vegetation | |||
+| Syn ICG Drone | ICG Drone | SkyScenes | SynDrone | VALID | SynthAer |
+| 1 other | obstacle | unlabelled | unlabelled | background | sky |
+| dog | other | other | animal | ||
+| conflicting | traffic sign | traffic sign | sign | ||
+| ar-marker | sky | sky | ice | ||
+| unlabelled | bridge | bridge | bridge | ||
+| gurad rail | guard rail | stones | |||
+| traffic light | traffic light | traffic light | |||
+| static | static | pierruble | |||
+| dynamic | dynamic | garbage bin | |||
+| plane | |||||
+| harbor | |||||
+| land | |||||
+| chair | |||||
+| ship | |||||
+| lamp | |||||
+| tunnel | |||||
+| bus stop | |||||
+| powerline | |||||
+| garbage bin | |||||
+| other low obstacle | |||||
+| other high obstacle | |||||
+| 2 fence | fence | fence | fence | fence | |
+| 3 pole | fence pole | pole | pole | ||
+| 4 vegetation | tree | vegetation | vegetation | tree | tree |
+| bald-tree | other plant | vegetation | |||
+| 5 building | wall | wall | wall | wall | |
+| roof | building | building | building | ||
+| door | |||||
+| window | |||||
+| 6 water | water | water | water | water | |
+| pool | pool | ||||
+| 7 bicycle | bicycle | bicycle | bicycle | ||
+| motorcycle | motorcycle | ||||
+| 8 vehicle | car | car | car | small vehicle | vehicle |
+| truck | truck | large vehicle | |||
+| bus | bus | ||||
+| train | |||||
+| 9 person | person | pedestrian | pedestrian | person | |
+| rider | rider | ||||
+| 10 paved area | paved area | road | road | road | road |
+| sidewalk | sidewalk | pavement | footpath | ||
+| ground | ground | ||||
+| roadline | roadline | ||||
+| rail track | rail track | ||||
+| 11 terrain | rocks | terrain | terrain | land | |
+| gravel | |||||
+| dirt | |||||
+| vegetation | |||||
 | grass |
-SyntheticReal.
-As noted in Sec of the main paper, since different real-world datasets have different class vocabularies and definitions, for our SyntheticReal semantic segmentation experiments, we adapt the class-vocabulary of the synthetic source dataset (SkyScenes, SynDrone) to that of the target real dataset (UAVid, AeroScapes, ICG Drone).
+SyntheticReal. As noted in Sec. of the main paper, since different real-world datasets have different class vocabularies and definitions, for our Synthetic
+Real semantic segmentation experiments, we adapt the class-vocabulary of the synthetic source dataset (SkyScenes, SynDrone, VALID, SynthAer) to that of the target real dataset (UAVid, AeroScapes, ICG Drone).
 This is done using a class-merging scheme based on the class-vocabularies and after visually inspecting dataset annotations.
-We provide the class-merging schemes used for both SkyScenes and SynDrone [25] (synthetic) across real counterparts UAVid [19] (in Table.
-8), AeroScapes [22] (in Table.
+We provide the class-merging schemes used for the synthetic datasets (SkyScenes, SynDrone [42], VALID [5] and SynthAer [47]) across real counterparts UAVid [27] (in Table.
+7), AeroScapes [34] (in Table.
 9) and ICG Drone [20] (in Table.
 10).
+We also include a coarser common merging scheme for SkyScenes and all the real datasets UAVid, AeroScapes, and ICG Drone in Table.
+8.
 SkyScenes Diagnostic Experiments.
-To assess the sensitivity of trained models to different factors – weather, time of day, height, pitch, etc.– we train models on different SkyScenes variations and evaluate them on held-out-variations.
-For these experiments, we recude the SkyScenes vocabulary to a reasonable subset of classes (consistent with the widely used Cityscapes [5] palette) – road, sidewalk, building, wall, fence, pole, traffic light, traffic sign, vegetation, water, sky, pedestrian, rider, cars, truck, bus, roadline, motorcycle, bicycle and an ignore class.
-C.2 Training, Validation and Test Splits
-C.2.1 Synthetic Real Experiments
+To assess the sensitivity of trained models to different factors – weather, time of day, height, pitch, etc.– we train models on different SkyScenes variations and evaluate them on held-out-variati-ons.
+For these experiments, we reduce the SkyScenes vocabulary to a reasonable subset of classes (consistent with the widely used Cityscapes [8] palette) – road, sidewalk, building, wall, fence, pole, traffic light, traffic sign, vegetation, water, sky, pedestrian, rider, cars, truck, bus, roadline, motorcycle, bicycle and an ignore class.
+0.C.2 Training, Validation and Test Splits
 SkyScenes. For each
@@
 % ( images) each for validation and testing.
-SynDrone. SynDrone has images per town (across Carla towns) for each of the
-combinations,
-resulting in a total of images per
-combination.
-We use of these datapoints for training models, with kept aside for testing. The datapoints selected for training and testing are kept consistent with the one reported in SynDrone [25].
-C.2.2 SkyScenes Diagnostic Experiments
-Weather & Daytime Variations. For each weather variation we sample evenly across the towns and
-combinations (excluding ),
+While creating train, val and test splits, we collect equal number of samples from each town by dividing each town-specific traversal sequence into segments: the initial % for training, the next % for testing, and the final % of the segment for validation.
+Moreover, within each split, we ensure that every viewpoint is accompanied by its different variations across weather, daytime, height, and pitch settings.
+This safeguards against any potential cross-contamination across different splits while ensuring fair representation and equal distributions of all variations.
+SynDrone. SynDrone has images per town (across Carla towns) for each of the combinations,
+resulting in a total of images per combination.
+We use of these data points for training models, with kept aside for testing.
+The data points selected for training and testing are kept consistent with the one reported in SynDrone [42].
+VALID.VALID has a total of images spread across different height variations and different layout and daytime conditions all from the nadir perspective.
+For the combination, and , the total images are , and respectively which are split into for training and each for validation and testing.
+The data points selected for training and testing are kept consistent with the one reported in VALID [5]
+SynthAer. The split of the dataset and the data points selected for training and testing are consistent with the one reported in SynthAer [47] with images for training, for validation and for testing.
+0.C.3 SkyScenes Diagnostic Experiments
+Weather & Daytime Variations. For each weather variation we sample evenly across the towns and combinations (excluding ),
 resulting in a total of images.
-We use
-% ( images) of these
-for training models, and remaining % ( images) each for validation and testing.
+We use % ( images) of these for training models, and remaining % ( images) each for validation and testing.
 We evaluate the model on the same weather variation it was trained on and select the model with the best mIoU score for further evaluations on other weather variations.
@@
 For each variation, we evenly sample across towns and weather and daytime conditions, resulting in a total of images per height and pitch variation.
-Out of these, % ( images) are allocated for training models, and % ( images) each for validation and testing.We evaluate the model on the same height & pitch setting it was trained on and select the model with the best mIoU score for further evaluations on other height & pitch settings.
-C.3 Training Details
+Out of these, % ( images) are allocated for training models, and % ( images) each for validation and testing.
+We evaluate the model on the same height & pitch setting it was trained on and select the model with the best mIoU score for further evaluations on other height & pitch settings.
+0.C.4 Training Details
 Semantic Segmentation.
-For our semantic segmentation experiments we use both CNN – DeepLabv2 [3] (ResNet-101 [10] backbone) – and vision transformer – DAFormer [11] (with HRDA [12] source training; MiT-B5 [36] backbone) – based semantic segmentation architectures.
+For our semantic segmentation experiments we use both CNN – DeepLabv2 [4] (ResNet-101 [14] backbone) – and vision transformer – 1.
+DAFormer [16] (with HRDA [17] source training; MiT-B5 [55] backbone) and 2.
+Rein DINOv2 [54] – based semantic segmentation architectures.
 Note that we utilize only the DAFormer architecture to perform source-only training for our experiments.
 Following [18], we enable rare class sampling [16] and use Imagenet feature-distance for our thing classes during training.
-All models are trained using the AdamW [18] optimizer coupled with a polynomial learning rate scheduler with an initial learning rate of .
+DeepLabv2 [4] and DAFormer [16] are trained using the AdamW [26] optimizer coupled with a polynomial learning rate scheduler with an initial learning rate of .
+For our fine-tuning experiments, we use an initial learning rate of .
+For Rein [54], we use initial learning rate of and for fine-tuning we use .
 Each model is trained for k iterations with a batch size of .
-For our fine-tuning experiments, we use an initial learning rate of .
 Depth-Aided Semantic Segmentation.
@@
 We train both RGB and RGB+D models with complete supervision for (moderate viewpoint) conditions on SkyScenes.
-C.4 Evaluation Details
+0.C.5 Evaluation Details
 Due to memory constraints, in addition to heavily parameterized models, our GPUs were unable to fit images larger than .
@@
 Similarly, for ICG Drone, we obtain overall image prediction using such crop predictions.
-Appendix D Results
-D.1 SyntheticReal Aligned Data Selection
-As stated in the main paper, for our SyntheticReal experiments, we train models on subsets of synthetic datasets that are aligned with corresponding real data conditions.
-In case of UAVid and AeroScapes we find that viewpoints in SkyScenes and viewpoints in SynDrone are best aligned with UAVid conditions (see Table.
-11 and Table.
-12) and provide best transfer performance.
-Similarly, for ICG Drone we observe that SkyScenes conditions are best aligned with the low-altitude, nadir perspective imagery in ICG Drone and lead to best transfer performance (see Table.
-13).
-However, for SynDrone, we find that the model trained on has best transfer performance, indicating that model performance is more sensitive to pitch alignment than height alignment.
+Appendix 0.D Results
+0.D.1 Ground vs Aerial Performance Gap
+Recent advancements in synthetic-to-real (syn-to-real) generalization have significantly closed the performance gap in ground imagery to about 20%, unlike in aerial imagery where the gap remains as high as 50%.
+In our study, we evaluated Rein DINOv2 [54] across ground-view settings using GTAVCityscapes and aerial-view settings using SkyScenes ICG.
+For ground views, the gap between GTAVCityscapes and Cityscapes [8] mIoU was 16.29%, demonstrating a mIoU of compared to .
+Aerial views showed a larger gap, with SkyScenes ICG mIoU at versus ICG mIoU at , resulting in a 50.53% gap.
+This highlights the need for focused efforts to address the larger performance discrepancies observed in aerial imagery.
 | h(m) | Synthetic mIoU () | ||||||||
@@
 | h(m) | Synthetic mIoU () | |||||||||||||
-| other | fence | pole | vegetation | building | water | bicycle | vehicle | person | paved area | terrain | Avg | Avg | ||
+| other | fence | pole | vegetation | building | water | bicycle | vehicle | person | paved area | terrain | Avg | Avg∗ | ||
 | SkyScenes | ||||||||||||||
@@
 | 15 | ||||||||||||||
+0.D.2 Advantage of SkyScenes over Real Datasets
+Curating densely annotated real data under diverse conditions in a controlled / uncontrolled manner is prohibitively expensive.
+As a synthetic alternative, SkyScenes is well-suited for evaluating model sensitivity to changing conditions.
+Further, due to the lack of diverse annotated data, we observe that real aerial datasets are somewhat homogenous.
+Consequently, models trained on real data under specific conditions (unlike SkyScenes) struggle to generalize to differing conditions – Rein DINOv2 [54] models trained on SkyScenes generalize better to ICG Drone [20] ( mIoU) compared to ones trained on AeroScapes [34] ( mIoU) and UAVid [27] ( mIoU)666A common merging scheme was used as detailed in Table.
+8
+0.D.3 SyntheticReal Aligned Data Selection
+As stated in the main paper, for our SyntheticReal experiments, we train models on subsets of synthetic datasets that are aligned with corresponding real data conditions.
+In case of UAVid and AeroScapes we find that viewpoints in SkyScenes and viewpoints in SynDrone are best aligned with UAVid conditions (see Table.
+11 and Table.
+12) and provide best transfer performance.
+Similarly, for ICG Drone we observe that SkyScenes conditions are best aligned with the low-altitude, nadir perspective imagery in ICG Drone and lead to best transfer performance (see Table.
+13).
+However, for SynDrone, we find that the model trained on has best transfer performance, indicating that model performance is more sensitive to pitch alignment than height alignment.
+0.D.4 SyntheticReal Additional Experiments
+| (Source) | (Target) Real-World mIoU () | ||
+| Synthetic | UAVid | AeroScapes | ICG Drone |
+| 1 SynthAer | |||
+| 2 VALID | |||
+| 3 SynDrone | 54.92 | ||
+| 4 SkyScenes | 43.96 | 28.10 | |
+A proper comparison with other synthetic datasets on a recognition task would require similar classes, (which a subset of synthetic alternatives, such as MidAir [12], do not satisfy).
+Considering other desired attributes (outlined in Table.1 in the paper), we chose SynDrone [42] as the most aligned alternative for our experiments.
+For comprehensive analysis and additional comparisons with other synthetic alternatives across various SyntheticReal settings, we included comparisons with VALID [5], which features only nadir perspective, and SynthAer [47], which lacks human classes, in Table 14.
+From the results in Table 14 , it is evident that models trained on SkyScenes generalize better to real datasets compared to other synthetic alternatives.
+0.D.5 SkyScenes Real Data Experiments
 | Training | Data Size | Synthetic mIoU () | |||||||
@@
 | 20 JT | 70.01 | ||||||||
-| Target | |||||||||
-| 21 Target | |||||||||
-| 22 Target | |||||||||
-| 23 Target | |||||||||
-| 24 Target | |||||||||
-| 25 Target | |||||||||
-| Training | Data Size | Synthetic mIoU () | |||||||
-| Clutter | Building | Road | Tree | Low Vegetation | Human | Vehicle | Avg | ||
-| SkyScenes | |||||||||
-| 1 FT | |||||||||
-| 2 FT | |||||||||
-| 3 FT | |||||||||
-| 4 FT | |||||||||
-| 5 FT | |||||||||
-| 6 JT | 62.97 | ||||||||
-| 7 JT | 67.58 | ||||||||
-| 8 JT | 70.20 | ||||||||
-| 9 JT | 71.83 | ||||||||
-| 10 JT | |||||||||
-| SynDrone | |||||||||
-| 11 FT | |||||||||
-| 12 FT | |||||||||
-| 13 FT | |||||||||
-| 14 FT | |||||||||
-| 15 FT | |||||||||
-| 16 JT | |||||||||
-| 17 JT | |||||||||
-| 18 JT | |||||||||
-| 19 JT | |||||||||
-| 20 JT | |||||||||
 | Target | |||||||||
@@
 | 25 Target | ||||||||||
+| Training | Data Size | Synthetic mIoU () | |||||||
+| Clutter | Building | Road | Tree | Low Vegetation | Human | Vehicle | Avg | ||
+| SkyScenes | |||||||||
+| 1 FT | |||||||||
+| 2 FT | |||||||||
+| 3 FT | |||||||||
+| 4 FT | |||||||||
+| 5 FT | |||||||||
+| 6 JT | 62.97 | ||||||||
+| 7 JT | 67.58 | ||||||||
+| 8 JT | 70.20 | ||||||||
+| 9 JT | 71.83 | ||||||||
+| 10 JT | |||||||||
+| SynDrone | |||||||||
+| 11 FT | |||||||||
+| 12 FT | |||||||||
+| 13 FT | |||||||||
+| 14 FT | |||||||||
+| 15 FT | |||||||||
+| 16 JT | |||||||||
+| 17 JT | |||||||||
+| 18 JT | |||||||||
+| 19 JT | |||||||||
+| 20 JT | |||||||||
+| Target | |||||||||
+| 21 Target | |||||||||
+| 22 Target | |||||||||
+| 23 Target | |||||||||
+| 24 Target | |||||||||
+| 25 Target | |||||||||
 | Training | Data Size | Synthetic mIoU () | ||||||||
@@
 | 25 Target | ||||||||||
-D.2 SkyScenes Real Data Experiments
+| Training | Data Size | Synthetic mIoU () | |||||||
+| Clutter | Building | Road | Tree | Low Vegetation | Human | Vehicle | Avg | ||
+| SkyScenes | |||||||||
+| 1 FT | 70.7 | ||||||||
+| 2 FT | 73.99 | ||||||||
+| 3 FT | |||||||||
+| 4 FT | 76.44 | ||||||||
+| 5 FT | 76.89 | ||||||||
+| 6 JT | |||||||||
+| 7 JT | |||||||||
+| 8 JT | 75.08 | ||||||||
+| 9 JT | |||||||||
+| 10 JT | |||||||||
+| Target | |||||||||
+| 21 Target | |||||||||
+| 22 Target | |||||||||
+| 23 Target | |||||||||
+| 24 Target | |||||||||
+| 25 Target | |||||||||
 In addition to zero-shot transfer to real data, we also show how SkyScenes is useful as additional training data when labeled real-world data is available.
@@
 10 and Fig.
-11 we make a similar comparison with the DAFormer [11] architecture.
+11 we make a similar comparison with the DAFormer [16] architecture, and in Fig.
+12 with the Rein DINOv2 [54] architecture.
 In low-shot regimes (when little “real” world data is available), SkyScenes data (either explicitly via joint training or implicitly via finetuning) is beneficial in improving recognition performance.
@@
 In Table.
-13(a) and Table.
-13(c) we present similar fine-grained (per-class) comparison of SkyScenes with SynDrone for a DeepLabv2 model when real data is available for training – via Target-Only, Finetuning or Joint-Training.
-Tables.
-13(b) and 13(d) show similar comparisons using the DAFormer [11] model.
-D.3 SkyScenes Diagnostic Experiments
-Similar to Table 6 (d) in the main paper, in Table 15, we assess broader sensitivity of models by training DAFormer [11] models across all (total ) settings and evaluate them across the same conditions in SkyScenes.
-Table 15 (a) models trained on are representative of one extreme of the range (both lowest height and pitch values) – we notice that this model is extremely sensitive to height variations.
-From Tables 15 (a), (e) and (i), we can deduce that due to the high variability in perspective between and other conditions, models trained on do not generalize well to other values.
-On the other hand, Tables.
-15 (b), (c), (d), (f), (g) and (h) models trained on oblique perspectives are better at generalizing to other pitch conditions.
+15 and Table.
+16 we present similar fine-grained (per-class) comparison of SkyScenes with SynDrone for a DeepLabv2 model when real data is available for training – via Target-Only, Finetuning or Joint-Training.Tables.
+17 and 18 show similar comparisons using the DAFormer [16] model.
+Table.
+19 shows similar comparison using the Rein DINOv2 [54] model.
+0.D.6 SkyScenes Diagnostic Experiments
 | Test mIoU () | ||||
@@
 | 3 m | 29.58 | 42.59 | 50.73 | 57.07 |
-D.4 Qualitative examples
-In Fig.
-12, 13 and 14, we show qualitative examples of predictions made by SkyScenes and SynDrone trained DeepLabv2 and DAFormer models on UAVid, AeroScapes, and ICG Drone respectively.
-In Fig.
-15 and 16, we show how predictions are impacted by changing SkyScenes conditions.
+Similar to Table 6 (d) in the main paper, in Table 20, we assess broader sensitivity of models by training DAFormer [16] models across all (total ) settings and evaluate them across the same conditions in SkyScenes.Table 20 (a) models trained on are representative of one extreme of the range (both lowest height and pitch values) – we notice that this model is extremely sensitive to height variations.Tables 20 (a), (e) and (i), we can deduce that due to the high variability in perspective between and other conditions, models trained on do not generalize well to other values.
+On the other hand, Tables.
+20 (b), (c), (d), (f), (g) and (h) models trained on oblique perspectives are better at generalizing to other pitch conditions.
+| Daytime | Test mIoU () | ||||
+| Conditions | Vanilla | Motion Blur | Spatter | ||
+| sev. = 3 | sev. = 5 | sev. = 3 | sev. = 5 | ||
+| Noon | 83.54 | 74.31 (9.23) | 67.49 (16.05) | 78.42 (5.11) | 63.03 (20.51) |
+| Sunset | 84.32 | 74.62 (9.70) | 68.08 (16.24) | 78.10 (6.22) | 68.62 (15.69) |
+0.D.7 Visual Artifacts
+Real-world images can have sensory artifacts beyond weather/daytime variations (UAV123 [30], VIRAT [35], UG2 [52]).
+Prior work (RobustNav [2]) shows it’s possible to simulate such corruptions with varying severities.
+Fig.0.D.6 provides examples of motion blur and spatter, highlighting how models trained on clean data degrade under these conditions (see Table.
+0.D.6).
+0.D.8 Auto labeling for Real Datasets
+Fig.13 shows a sample semantic segmentation prediction of a random UAVid image using Grounded-SAM [41].
+We find that most of the roads are absent from the prediction, no humans are predicted, and only a few instances of vehicles, buildings, and trees are identified correctly.
+While auto-labeling using foundation models has potential, we believe there is still room for improvement in specific applications (such as aerial).
+0.D.9 Qualitative examples
+Appendix 0.E Limitations
+Depth of distant objects in SkyScenes is imperfect, akin to real-world conditions.
+As evident from Fig.
+5 in the paper, depth observations struggle to discern finer details of high-altitude viewpoints.
+We hypothesize that improvement from depth is helpful only for objects close to the camera or lower altitude viewpoints.
+Appendix 0.F Future Work
+We plan on updating SkyScenes with evolving considerations for real-world aerial scene-understanding – improved realism, additional anticipated edge cases – as more and more features are supported in the underlying simulator and provide additional support for 3D perception tasks (see Fig.
+19.)